### v1

# SAQR-QC: A Logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits

Abstract.
Reasoning about quantum programs remains a fundamental challenge, regardless of the programming model or computational paradigm. Despite extensive research, existing verification techniques are insufficient—even for quantum circuits, a deliberately restricted model that lacks classical control, but still underpins many current quantum algorithms. Many existing formal methods require exponential time and space to represent and manipulate (representations of) assertions and judgments, making them impractical for quantum circuits with many qubits. This paper presents a logic for reasoning in such settings, called SAQR-QC. The logic supports Scalable but Approximate Quantitative Reasoning about Quantum Circuits, whence the name. SAQR-QC has three characteristics: (i) some (deliberate) loss of precision is built into it; (ii) it has a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small; and (iii) most importantly, to make reasoning scalable, all reasoning steps are local—i.e., they each involve just a small number of qubits. We demonstrate the effectiveness of SAQR-QC via two case studies: the verification of GHZ circuits involving non-Clifford gates, and the analysis of quantum phase estimation—a core subroutine in Shor’s factoring algorithm.
copyright: nonejournal: PACMPLjournalvolume: 1journalnumber: CONFarticle: 1journalyear: 2018publicationmonth: 1
1. Introduction
Quantum computing, as a new computational paradigm, exploits quantum parallelism and interference to outperform classical methods on specific tasks. For instance, Shor’s algorithm (Shor94, ) efficiently factors large integers using the Quantum Fourier Transform (QFT). Kitaev’s phase estimation (Kitaev1995, ) serves as a foundation for quantum simulation (Aspuru-Guzik2005, ) and optimization algorithms, such as QAOA (FarhiGoldstoneGutmann14, ). Interestingly, key components that provide the advantages of quantum computation over classical computation are obtained (1) without the use of classical control-flow constructs such as while loops, and (2) with a fixed number of qubits, the quantum analogue of bits, throughout the computation.
Despite quantum computing being a much different computational paradigm, one is still interested in showing that programs behave as desired. Significant effort has been put into developing methods to verify quantum programs using classical computers (BHY19, ; ZYY19, ; zhou2021quantum, ; 10.1145/3453483.3454029, ; Unruh, ; YP21, ; DBLP:journals/pacmpl/0001CLLTY23, ; 10.1145/3704873, ; 10.1145/3519939.3523431, ). Particular noteworthy is Ying’s Quantum Hoare Logic (QHL) (Ying11, ), which extends classical Hoare logic to reason about the partial correctness of quantum programs, using quantum predicates, semidefinite positive Hermitian operators, as preconditions and postconditions. The quantum Hoare triple, introduced in (DP06, ), is defined as
(1) $\{A\}P\{B\}\quad\mathrm{iff}\quad\textrm{for all}~{}\rho,{\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\llbracket P\rrbracket(\rho)),$
where $A$ and $B$ are quantum observables representing preconditions and postconditions (and ${\mathrm{Tr}}$ is the matrix-trace operation).
Despite this extensive body of research, quantum program verification remains inadequate, even for quantum circuits—a deliberately restricted model that lacks classical control in which many current quantum algorithms are formulated. Existing formal methods struggle to scale; for instance, although QHL provides a comprehensive framework for tracking quantum program behavior, including success probabilities, it lacks support for local reasoning and requires exponential time and space, making it impractical for large systems. More precisely, for a unitary program $P$, whose semantics corresponds to a unitary matrix $U_{P}$, QHL provides the rule
(2) $\{U_{P}^{{\dagger}}BU_{P}\}P\{B\}$
Although this rule is complete for unitary programs, it imposes a computationally intractable task on the verification process—namely, computing $U_{P}^{\dagger}BU_{P}$ for a general unitary $U_{P}$. Unfortunately, even representing a general $n$-qubit unitary matrix (or a general observable) requires $(2^{n})^{2}=4^{n}$ parameters.
This discussion raises the following question: {mdframed} Is there a theoretical foundation for scalable quantitative reasoning about quantum programs that use many qubits? Here, “scalable” means that the computational resources required for verification grow polynomially with the number of qubits.
This paper answers this question in the affirmative. It presents a logic for reasoning about programs expressed as quantum circuits, called SAQR-QC—for Scalable but Approximate Quantitative Reasoning about Quantum Circuits. SAQR-QC has three characteristics: (i) some (deliberate) loss of precision is built into it; (ii) it has a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small; and (iii) most importantly, to make reasoning scalable, all reasoning steps are local—i.e., they each involve just a small number of qubits.
SAQR-QC applies to the circuit model of quantum programs, which is widely used in quantum computing. Most known quantum algorithms either fit within the quantum-circuit model or consist of some number of runs of a quantum circuit. We consider quantum circuits that act on $n$ qubits and are composed of a polynomial number of unitary gates of up to $k$ qubits (where $k$ is a constant independent of $n$).111 We place no restrictions on the gate set, except that each elementary gate operates on a constant number of qubits. While our examples use one- and two-qubit gates, our results extend to circuits composed of gates acting on three qubits (such as the Toffoli gate), four qubits, etc. With quantum circuits, the input state is typically initialized to some computational basis state, such as |$\ket{0}^{\otimes n}$.
Does this setting —-quantum circuits composed of a polynomial number of $k$-qubit gates acting on |$\ket{0}^{\otimes n}$ —-make the problem simple enough to be efficiently solved by a classical computer? Not at all. Even if we only care about the value of the output state’s first qubit, which can be represented by a $2\times 2$ density matrix, and even if we restrict our interest to just the first diagonal entry of that matrix, the problem remains hard. Determining whether this single number is smaller than $1/3$ or larger than $2/3$ is believed not to be solvable by any classical polynomial-time algorithm unless quantum computation offers no advantage over classical computation.
One of the inspirations for our work is Quantum Abstract Interpretation (QAI) (YP21, ), a framework proposed by Yu and Palsberg that extends classical abstract interpretation (DBLP:conf/popl/CousotC77, ) to reason about quantum circuits. A QAI judgment for a quantum circuit $C$ is formalized as follows:
(3) $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}\quad\mathrm{iff}\quad\textrm{for all}~{}\rho,\rho\vDash\mathcal{P}~{}\mathrm{implies}~{}\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},$
where $\mathcal{P},\mathcal{Q}$ are tuples of local projections used as abstract states. The above triple asserts that if the program’s input satisfies $\mathcal{P}$, the output will satisfy $\mathcal{Q}$, where satisfaction means that the output state lies within the space defined by the intersection of the spaces of the local projections in $\mathcal{Q}$.
QAI facilitates the analysis and verification of quantum programs by (a) mapping quantum states to an abstract domain that capture essential properties of the quantum state, and (b) providing a technique for applying abstract transformers in which all reasoning is local—i.e., each application of an abstract transformer involves just a small number of qubits. On the plus side, QAI enables polynomial-time analysis. However, on the minus side, QAI does not support the reasoning of quantitative. The local-reasoning method in QAI focuses on local projections, which provide qualitative insight but fail to capture success probabilities. However, most quantum algorithms succeed with probability less than 1, underscoring the need for a scalable framework for quantitative analysis of quantum programs—a challenge that remains unaddressed. Here, scalability requires verification algorithms that are significantly more efficient in time and space than brute-force simulation, which scales exponentially with the number of qubits.
QAI and QHL stand on opposite ends of a feature spectrum: whereas QAI is scalable but lacks quantitative reasoning, QHL offers quantitative reasoning, but is not scalable. SAQR-QC takes inspiration from both. Integrates QHL-like quantitative reasoning with the scalability of QAI to achieve scalable quantitative reasoning about quantum circuits with |$\ket{0}^{\otimes n}$ as input state.
In SAQR-QC, the idea behind the QHL-like component is to replace a single observable with a tuple of local observables, each represented by a semidefinite matrix that captures information about the corresponding reduced density matrix (NI11, ).
Notably, a tuple of reduced density matrices, each defined on a constant number of qubits, requires only a linear number of parameters to describe, contrasting sharply with the exponential complexity needed to represent the full quantum state. This locality enables tractable analysis and verification techniques, making it feasible to reason about large-scale quantum computations through their low-dimensional marginals. A judgment in SAQR-QC takes the following form:
(4) $\{\mathscr{A}|\mathcal{P}\}C\{\mathscr{B}|\mathcal{Q}\}.$
Here, $\mathscr{A}$ and $\mathscr{B}$ are tuples of local observables used to track the probability of success, where locality means that each observable acts non-trivially on only a constant number of qubits, independent of the total number of qubits. This locality ensures the potential for scalable reasoning in a manner similar to that of QAI. Likewise, $\mathcal{P}$ and $\mathcal{Q}$ are tuples of local projections, as defined in QAI. 222 The two-part structure of each assertion is analogous to how formulas in reasoning systems based on separation logic have two parts—a spatial formula and a pure logical formula) (DBLP:conf/fmco/BerdineCO05, ; DBLP:conf/tacas/DistefanoOY06, ). In both cases, the principle is to carry around two different kinds of constraints on the state. In SAQR-QC, the QAI-like component provides spatial information about the program state—i.e., the quantum state lies within a restricted subspace. The local observables of the QHL-like component then provide a bound on the inner product between the program output and a local observable. More precisely, the triple in Equation 4 asserts
(5) $\textrm{for all}~{}\rho,\rho\vDash\mathcal{P}~{}\mathrm{implies}~{}(i)~{}\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},~{}~{}\mathrm{and}~{}~{}(ii)~{}{\mathrm{Tr}}(M_{\mathscr{A}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),$
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are derived from $\mathscr{A}$ and $\mathscr{B}$, with the precise definition provided in § 4.
SAQR-QC has built into it some (deliberate) loss of precision, but the use of mechanisms from both QHL and QAI causes the accumulated loss of precision during a sequence of reasoning steps to remain small. Both the deliberate-loss-of-precision aspect and the use of two interacting formalisms are inspired by abstract interpretation.
An important use-case that we envision for SAQR-QC is for forward reasoning—i.e., given the pre-state assertion $\{\mathscr{A}|\mathcal{P}\}$333 Even though we restrict attention to circuits with input |$\ket{0}^{\otimes n}$, stating the pre-state assertion $\{\mathscr{A}|\mathcal{P}\}$ is not necessarily trivial. For the $\mathscr{A}$ component, one needs to choose what local observables to track; For the $\mathcal{P}$ component, one needs to choose what local projections to track. and circuit $C$, find a suitable post-state assertion $\{\mathscr{B}|\mathcal{Q}\}$. Obviously, the post-state assertion obtained using SAQR-QC will not characterize the strongest post-condition of $C$ with respect to $\mathscr{A}|\mathcal{P}$. Instead, the goal is that the obtained post-condition will be strong enough to obtain useful information about the computation.
One important instance of such information is the computation’s success probability, which, in models like BQP,444 “BQP is the class of decision problems solvable by a quantum computer in polynomial time with bounded error,” just as “P is the class of decision problems solvable by a classical deterministic computer in polynomial time.” is typically determined by a diagonal entry of the output state’s first qubit.
It should be borne in mind that SAQR-QC is a logic, not an automated reasoning system. SAQR-QC is similar to other logics in that clever insights and eureka steps are sometimes needed to be able to establish that certain properties do, in fact, hold. One of our goals is to be able to establish success probabilities for quantum circuits, and doing so with SAQR-QC is not “cookbook.” For instance, the two case studies presented in § 5 and § 6 are both examples in which the specific choices of the predicates used are crucial to obtaining the sharp results detailed in those sections.
Contributions
The work described in the paper makes the following contributions:
  • We present SAQR-QC, which supports—and is named for—Scalable but Approximate Quantitative Reasoning about Quantum Circuits.
  • SAQR-QC takes inspiration from QHL: similar to QHL, it supports quantitative reasoning.
  • SAQR-QC takes inspiration from QAI: similar to QAI, some (deliberate) loss of precision is built into the reasoning rules of SAQR-QC. The advantage is the same as in QAI, namely that because all reasoning steps are local—i.e., they each involve just a small number of qubits—they can be performed in polynomial time, and thus the reasoning method is scalable.
  • SAQR-QC takes inspiration from abstract interpretation: in particular, the use of the QAI-like information to restrict the QHL-like information is a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small, and hence similar to a reduced product (DBLP:conf/popl/CousotC79, , §10.1).
  • We demonstrate the effectiveness of SAQR-QC via two case studies:
    • The verification of GHZ circuits involving non-Clifford gates: We apply SAQR-QC to derive a judgment that gives a characterization of the output state that is precise in every aspect except for a phase factor.
    • The analysis of quantum phase estimation—a core subroutine in Shor’s factoring algorithm: We apply SAQR-QC to derive a judgment showing that, for a given constant $k$, the algorithm provides the best estimation of the last $k$ bits of the phase with probability at least $\frac{4}{\pi^{2}}$. To obtain this result, we first develop a lossless local-reasoning method for the quantum Fourier transform using QAI. In particular, the concretization of the abstract output state exactly matches the true output state for computational-basis inputs. To the best of our knowledge, no prior approach has shown that within polynomial time it is possible to obtain such results about either the quantum Fourier transform or quantum phase estimation.
Organization of the paper
We present background material about quantum computing in § 2. § 3 summarizes the basics of QAI (YP21, ). § 4 defines local observables as predicates, and formalizes the judgments of SAQR-QC, which integrate local observables and QAI. We employ SAQR-QC to reason about the general GHZ circuit (§ 5) and quantum phase estimation (§ 6). § 7 discusses related work. § 8 concludes. Some proofs and derivations are available in the paper’s appendices, which are submitted as Supplementary Material.
2. Background & Notation
To make the paper self-contained, this section provides background material, and discusses notation for—and properties of—quantum computing. The material in this section is similar to what can be found in published books and papers, e.g., (NI11, ; YP21, ; Ying11, ). Readers already familiar with quantum computing may wish to proceed directly to § 3.
2.1. Preliminaries
We use the notation $[n]$ to denote the set $\{1,\ldots,n\}$, $\setminus$ to denote set difference, and $|s|$ to denote the cardinality of a set $s$. We assume familiarity with Dirac notation and standard linear algebra concepts, including Hilbert space, tensor products, orthonormal bases, and inner and outer products.
Linear operators on $d$-dimensional complex vector spaces are represented as $d\times d$ matrices over $\mathbb{C}$, denoted $\mathbb{C}^{d\times d}$. The identity operator is denoted by $I$. For an operator $A$, the conjugate transpose is defined as $A^{\dagger}=(A^{T})^{*}$, where $A^{T}$ is the transpose and $(\cdot)^{*}$ denotes complex conjugation. An operator $A$ is Hermitian if $A=A^{\dagger}$, and positive semi-definite if all its eigenvalues are non-negative. The trace of a matrix $A$, denoted ${\mathrm{Tr}}(A)$, is the sum of its diagonal entries: ${\mathrm{Tr}}(A)=\sum_{i}A_{ii}$.
The Löwner order, defined on Hermitian matrices in quantum mechanics and convex analysis, establishes a relationship $A\leq B$ indicating that $B-A$ is positive semidefinite. The Löwner order is vital for comparing states and operators.
2.2. Quantum States
A quantum state describes the state of a quantum system. For a single qubit, the state |$\ket{\psi}$ belongs to a two-dimensional Hilbert space and can exist as a superposition of the basis states |$\ket{0}$ and |$\ket{1}$. For an $n$-qubit system, the state resides in a $2^{n}$-dimensional Hilbert space and can exhibit both complex superpositions and entanglement among qubits.
Quantum systems may also be in mixed states, represented by a density matrix $\rho$, which generalizes pure states to account for probabilistic mixtures—capturing both classical and quantum uncertainty. Quantum operations, whether on pure or mixed states, include unitary transformations (which preserve total probability) and measurements, which probabilistically collapse the system to a classical outcome based on the state’s amplitudes.
2.3. Reduced Density Matrices
Let $\mathbb{C}^{d_{1}}$ and $\mathbb{C}^{d_{2}}$ be the Hilbert spaces of two quantum systems. The state space of the composite system is modeled by the tensor product $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}$, and analyzing subsystems requires the notion of the partial trace. Formally, the partial trace over $\mathbb{C}^{d_{1}}$, denoted ${\mathrm{Tr}}_{1}(\cdot)$, maps operators on $\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}$ to operators on $\mathbb{C}^{d_{2}}$, and is defined by
[…]
for all |$\ket{\varphi_{1}},\ket{\psi_{1}}\in\mathbb{C}^{d_{1}}$ and |$\ket{\varphi_{2}},\ket{\psi_{2}}\in\mathbb{C}^{d_{2}}$, and extended linearly. Similarly, ${\mathrm{Tr}}_{2}(\cdot)$ denotes the partial trace over $\mathbb{C}^{d_{2}}$. For a composite system with density matrix $\rho\in\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}}$, the reduced states of the first and second subsystems are given by ${\mathrm{Tr}}_{2}(\rho)$ and ${\mathrm{Tr}}_{1}(\rho)$, respectively.
The notion of partial trace extends naturally to $n$-partite systems: for a subset $s\subseteq[n]$, the reduced density matrix of the subsystem indexed by $s$ is given by
$\rho_{s}={\mathrm{Tr}}_{[n]\setminus s}(\rho),$
where ${\mathrm{Tr}}_{[n]\setminus s}$ denotes tracing out all qubits not in $s$. The partial-trace operation preserves positive semi-definiteness (NI11, ).
Reduced density matrices play a central role in the analysis of multipartite quantum systems. Many important properties of a quantum system depend solely on the reduced density matrices of its subsystems. This principle holds in quantum computation as well. For instance, the success probability of many quantum algorithms—such as the well-known HHL algorithm (HHL08, )—depends solely on the reduced density matrix of a specific qubit (the “signal” qubit). In such algorithms, the signal qubit is measured, and the computation is considered successful if the outcome is |$\ket{1}$. Importantly, the probability of this outcome is fully determined by the reduced density matrix of the signal qubit, regardless of the entanglement or structure of the global quantum state.
2.4. Unitary Operations
Unitary operations are fundamental transformations in quantum mechanics; they preserve the norm of a quantum state and are represented by a unitary matrix. A unitary matrix $U$ satisfies $U^{\dagger}U=I$, where $U^{\dagger}$ denotes the conjugate transpose of $U$ and $I$ is the identity matrix. These operations are crucial for manipulating quantum states and implementing quantum algorithms. For a pure state |$\ket{\psi}$, applying a unitary operator $U$ transforms |$\ket{\psi}$ to $U\ket{\psi}$. For a density operator $\rho$, the transformation is $\rho\mapsto U\rho U^{\dagger}$.
Commonly used single-qubit operators include the Hadamard gate $H$, the $T$ gate, the family of gates $\set{R_{m}\mid m\in\mathbb{N}}$ and the Pauli gates $I$, $X$, $Y$, and $Z$. Commonly used two-qubit gates include the SWAP operation SWAP and the controlled-NOT operation CNOT.
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\qquad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\qquad Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\qquad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\qquad\textit{SWAP}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix}$
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\qquad T=\begin{pmatrix}1&0\\ 0&e^{i\pi/4}\end{pmatrix}\qquad R_{m}={\begin{pmatrix}1&0\\ 0&e^{2\pi i/2^{m}}\end{pmatrix}}\qquad\textit{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$
2.5. Observables
Quantum observables are physical quantities in a quantum system that can be measured, such as position, momentum, energy, and spin. Mathematically, observables are represented by Hermitian (self-adjoint) operators on a Hilbert space, satisfying $O^{\dagger}=O$, which ensures that measurement outcomes are real. In certain contexts—such as Quantum Hoare Logic (QHL)—it is common to further restrict observables to the range $0\leq O\leq I$, where the inequality is understood in the Löwner partial order (§ 2.1), meaning that both $O$ and $I-O$ are positive semidefinite. These bounded observables often serve as predicates or quantum effects, capturing partial truth values within the quantum program-verification framework.
2.6. Quantum Circuits and Semantics
In this work, we focus on quantum programs—represented as quantum circuits—operating on a fixed number of qubits, says $n$. A quantum program is composed of a sequence of unitary instructions $U_{F_{1}},\dots,U_{F_{|p|}}$. Each gate $U_{F_{\ell}}$ operates on a subset of qubits $F_{\ell}\subseteq[n]$. The initial state is taken to be |$\ket{0}^{\otimes n}=\ket{0^{n}}$, and the meaning (or semantics) of the program is given by the matrix product
$U_{F_{|p|}}\cdots U_{F_{1}}\ket{0^{n}}.$
To interpret $U_{F_{\ell}}$ as an $n$-qubit unitary, we embed it into the full register by tensoring with identity operators. If $F_{\ell}=\{i\}$, then the corresponding lifted unitary is
$U\otimes I_{[n]\setminus\{i\}}:=(\otimes_{k>i}I)\otimes U\otimes(\otimes_{0\leq j<i}I),$
where $U$ acts nontrivially on qubit $i$, and $I$ is the identity matrix on one qubit. We will explicitly subscript identity matrices to indicate which qubits they act upon.
For a two-qubit unitary $U_{F_{\ell}}$ acting on qubits $F_{\ell}=\{i,j\}$, we similarly interpret it as
$U\otimes I_{[n]\setminus\{i,j\}},$
with the appropriate placement determined by the positions of $i$ and $j$ in the register.
To simplify the presentation and reduce notational overhead, we formulate the semantics of quantum circuits assuming 2-qubit gates. Nevertheless, all of our methods naturally extend to circuits composed of gates acting on up to $m$ qubits, for any constant $m$.
Definition 2.1 (Syntax).
The syntax of quantum programs is given by
$C::=\mathbf{skip}\mid\bar{q}:=U[\bar{q}]\mid C_{1};C_{2}$
We write $\llbracket C\rrbracket$ to denote the semantics of a quantum program $C$. If $C$ represents a unitary transformation $U_{C}$, then for any input density matrix $\rho$, its semantics is given by
$\llbracket C\rrbracket(\rho):=U_{C}\rho U_{C}^{\dagger}.$
We also define the dual action on observables as:
$\llbracket C\rrbracket^{*}(A):=U_{C}^{\dagger}AU_{C},$
for any observable $A$.
2.7. Projections
An orthogonal projection matrix $P$ satisfies $P=P^{\dagger}=P^{2}$, which a stricter condition than the classical $P=P^{2}$. For short, we refer to such matrices as “projections.” For example, $|00\rangle\langle 00|+|11\rangle\langle 11|$ is a rank-2 projection that projects any 4-dimensional vector onto a 2-dimensional subspace.
Each projection $P$ corresponds to a unique subspace $S_{P}=\{v\mid Pv=v\}$, and we use the terms “projections” and “subspaces” interchangeably. This correspondence establishes a partial order: for projections $P$ and $Q$, we have $P\subseteq Q$ iff $S_{P}\subseteq S_{Q}$.
Projections are positive semi-definite. The support of a positive semi-definite matrix $A$, $\operatorname{supp}(A)$, is the subspace spanned by eigenvectors with nonzero eigenvalues. According to Birkhoff and von Neumann (BvN36, ), a density matrix $\rho$ satisfies a projection $P$, denoted by $\rho\vDash P$, if $\operatorname{supp}(\rho)\subseteq P$. This property is equivalent to $P\rho=\rho$.
2.8. Lemmas
Our development of SAQR-QC relies on three fundamental operations on operators:
  • Löwner order of operators (denoted $A\leq B$),
  • Partial trace and trace operators (denoted ${\mathrm{Tr}}_{s}$ and ${\mathrm{Tr}}$, where ${\mathrm{Tr}}_{s}$ traces out subsystem $s$),
  • Expansion of an operator via tensor product (denoted $A_{s}\otimes I_{[n]\setminus s}$, where the operator $A_{s}$ acts on subsystem $s$ and is expanded to the full system).
The following lemmas capture algebraic relationships among these operations. These lemmas are instrumental in establishing the correctness of our SAQR-QC framework in a purely algebraic style. In what follows, we assume $s\subseteq[n]$; that $P$ is a projection operator on $n$-qubit systems; that $A,B$ is a positive semidefinite matrix on an $n$-qubit space, $E$ be a matrix; and that $\rho$ is a quantum state.
Lemma 2.1.
Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$,
${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right).$
Lemma 2.2.
For two square matrices $B$ and $E$, ${\mathrm{Tr}}(BE)={\mathrm{Tr}}(EB)$.
Lemma 2.3.
For $A,B\geq 0$, ${\mathrm{Tr}}(AB)\geq 0$.
Lemma 2.4.
If $A\leq B$, then ${\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\rho)$ for any density operator $\rho$. In particular, ${\mathrm{Tr}}(A)\leq{\mathrm{Tr}}(B)$.
Lemma 2.5.
For a density matrix $\rho$ and a projection $P$, we have
${\mathrm{Tr}}(P\rho)=1\quad\Leftrightarrow\quad\operatorname{supp}(\rho)\subseteq P\quad\Leftrightarrow\quad\rho\vDash P.$
The proof of 2.1 is given in Appendix A. The other lemmas can be proven using the definition of trace, partial trace, and support.
3. Qualitative Predicates for Local Reasoning: Quantum Abstract Interpretation (YP21, )
Qualitative predicates serve as logical assertions about quantum states, such as whether a subsystem lies in a given subspace. When expressed as tuples of local projectors, they enable scalable local reasoning, inferring global behavior from partial system views. Quantum Abstract Interpretation (QAI) (YP21, ) systematically propagates such predicates through quantum circuits using projectors on small subsystems, avoiding the exponential cost of full-state analysis.
This section reviews local projective predicates, formalizes their semantics, and describes their transformation under unitaries. Soundness is established via support-based semantics and partial trace, forming a foundation for abstract reasoning in quantum programs.
Definition 3.1 ((YP21, )).
A tuple $(P_{s_{1}},\cdots,P_{s_{m}})$ is called a projective predicate if each $P_{s_{i}}$ is a projection, i.e., $P_{s_{i}}^{2}=P_{s_{i}}$. We use $\mathcal{P}$ (or $\mathcal{Q},\mathcal{R}$) to denote projective predicates. In particular, we write $\mathcal{I}:=(I_{s_{1}},\cdots,I_{s_{m}})$ to represent the identity predicate.
Remark. Projective predicates are also referred to as abstract states.
Definition 3.2 ((YP21, )).
An $n$-qubit quantum state $\rho$ satisfies a projective predicate $\mathcal{P}=(P_{s_{1}},\cdots,P_{s_{m}})$, denoted by
$\rho\vDash^{QAI}\mathcal{P},$
if for all $1\leq i\leq m$, we have $P_{s_{i}}\rho_{s_{i}}=\rho_{s_{i}}$, i.e., $\rho_{s_{i}}\vDash P_{s_{i}}$. Equivalently, this property holds if $\rho\vDash\gamma(\mathcal{P})$, where
$\gamma(\mathcal{P}):=\bigcap_{i}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Given a quantum circuit $\mathbf{C}$ and a state $\rho\vDash\mathcal{P}$, the abstract-interpretation method presented in (YP21, ) constructs a predicate $\mathcal{Q}$ such that the post-state $\llbracket\mathbf{C}\rrbracket(\rho)$ satisfies $\mathcal{Q}$, denoted as
$\vDash^{QAI}\{\mathcal{P}\}{C}\{\mathcal{Q}\}$
Theorem 3.1 ((YP21, )).
Let $U_{F}$ be a unitary gate applied to the qubit set $s(F)$, and let $\mathcal{P}=(P_{s_{1}},\cdots,P_{s_{m}})$ be a projective predicate. For each $s_{i}$, define
[…]
Then
$\rho\vDash^{QAI}\mathcal{P}\quad\Rightarrow\quad U_{F}\rho U_{F}^{\dagger}\vDash^{QAI}U^{\sharp}(\mathcal{P}).$
Theorem 3.2 ((YP21, )).
Let $P=\mathrm{span}\{\ket{a_{1}a_{2}\cdots a_{n}},\ket{b_{1}b_{2}\cdots b_{n}}\}$, where the product states |$\ket{a_{i}}$ and |$\ket{b_{i}}$ are not parallel for every $i\in[n]$. Then
$P=\gamma(\mathcal{P}),$
where $\mathcal{P}=(P_{1,2},\ldots,P_{n-1,n})$ and each $P_{i,i+1}=\mathrm{span}\{\ket{a_{i}a_{i+1}},\ket{b_{i}b_{i+1}}\}.$
4. Correctness Formulas and the Logical System SAQR-QC
In this section, we first motivate the use of local observables in the definition of quantum predicates. We then apply this idea to Quantum Hoare Logic (QHL) as an initial attempt to define a corresponding judgment. Through a simple example, we demonstrate that this approach can lead to a significant loss of precision in reasoning. To address this limitation, we subsequently integrate the QAI technique into the framework, leading to the formal definition of the judgments used in SAQR-QC and a systematic presentation of the inference rules of SAQR-QC.
4.1. Motivation.
In the analysis and verification of quantum programs, reasoning about the full quantum state quickly becomes infeasible due to the exponential growth of the state space with the number of qubits. To address this challenge, we shift our focus to reduced density matrices, which capture the behavior of local subsystems. Notably, a tuple of reduced density matrices—each defined on a constant number of qubits—requires only a linear number of parameters to describe. This property stands in stark contrast to the exponential size of the global state and provides a practical pathway to tractable reasoning.
Let $(\rho_{s_{1}},\dots,\rho_{s_{m}})$ denote a tuple of reduced density matrices, each corresponding to a small subset $s_{i}\subseteq[n]$ of the full system. Inspired by quantum Hoare logic (QHL), where a global positive semi-definite matrix $A$ is used as a predicate to track changes in the global state $\rho$, we generalize this approach by employing linear functionals—namely, local observables—to track the evolution of each $\rho_{s_{i}}$. This approach allows us to perform quantitative reasoning over the program’s behavior through its low-dimensional marginals.
Local observables, which act nontrivially on only a few qubits, serve as efficient probes for monitoring how the reduced density matrices evolve during program execution. Because observable expectation values determine measurable quantities like success probabilities, fidelities, and entanglement, they form a natural bridge between the semantics of quantum programs and their operational outcomes.
By tracing how these observables evolve under unitary transformations and partial traces, we establish a scalable and compositional framework for reasoning about quantum programs. This approach enables precise quantitative analysis without requiring reconstruction of the global state, laying the groundwork for scalable verification of large-scale quantum computations through their locally observable structure.
4.2. Quantitative Local Reasoning via Generalized Predicates
This subsection defines predicates for quantitative local reasoning by generalizing from local projections to local observables. While qualitative reasoning (as in (YP21, )) relies on tuples of projections to characterize logical properties of subsystems, our approach extends this framework to encompass a broader class of observables—specifically, positive semidefinite operators bounded by the identity.
These generalized predicates enable us to track quantitative aspects of quantum programs, such as probabilities and expectation values, by monitoring the evolution of reduced density matrices over small, constant-size subsystems. This generalization forms the foundation for scalable, compositional reasoning about quantum computations through low-dimensional, information-preserving abstractions.
Definition 4.1.
Given an $n$-qubit Hilbert space, an integer $m$, and an $m$-tuple $S=(s_{1},\cdots,s_{m})$ with each $s_{i}\subseteq[n]$ indicating a subset of qubits, the set of predicates is defined as:
[…]
where each $A_{s_{i}}$ is a local observable acting nontrivially only on subsystem $s_{i}$. We usually use $\mathscr{A}$ (or $\mathscr{B}$, $\mathscr{D}$) to denote predicates. The integer $m$ is called the size of the predicate.
The idea is to define a quantum predicate for each $s_{i}$, represented by an observable $A_{s_{i}}$, focusing solely on the state of the quantum registers in $s_{i}$.
Definition 4.2.
For a predicate $\mathscr{A}=(A_{s_{1}},\cdots,A_{s_{m}})$, its matrix representation $M_{\mathscr{A}}$ is defined as:
$M_{\mathscr{A}}=\sum_{i=1}^{m}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Definition 4.3.
The domain of a predicate $(A_{s_{1}},\cdots,A_{s_{m}})$, denoted by $\mathrm{dom}(A_{s_{1}},\cdots,A_{s_{m}})$, is $(s_{1},\cdots,s_{m})$.
We will see that the above definition of predicates serves to track changes in a specified tuple of reduced density matrices by means of linear functionals, facilitating quantitative reasoning in a scalable and compositional manner. By Lemma 2.1, the trace ${\mathrm{Tr}}(M_{\mathscr{A}}\rho)$ depends only on the reduced density matrices $\rho_{s_{i}}$ of $\rho$:
(6) ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}\left(\sum_{i=1}^{m}(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho\right)=\sum_{i=1}^{m}{\mathrm{Tr}}[(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho]=\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}\rho_{s_{i}}).$
We are also interested in the order relation between predicates.
Definition 4.4.
Given two predicates $(A_{s_{1}},\cdots,A_{s_{m}})$ and $(A_{s_{1}}^{\prime},\cdots,A_{s_{m}}^{\prime})$ over the same domains, we define
$(A_{s_{1}},\cdots,A_{s_{m}})\sqsubseteq(A_{s_{1}}^{\prime},\cdots,A_{s_{m}}^{\prime})\quad\text{if and only if}\quad A_{s_{i}}\leq A_{s_{i}}^{\prime}\text{for all }i.$
Lemma 4.1.
The matrix representation of predicates is monotonic with respect to this ordering—i.e.,
$\mathscr{A}\sqsubseteq\mathscr{B}\quad\Rightarrow\quad M_{\mathscr{A}}\leq M_{\mathscr{B}}.$
4.3. First Attempt at Defining Judgments
This section presents a first attempt to introduce a framework for quantitative local reasoning in general quantum programs using tuples of local observables.
In Quantum Hoare Logic (QHL) (Ying11, ), the proof rule for a unitary operation $P$ takes the form
$\{U_{P}^{\dagger}BU_{P}\}\ P\ \{B\},$
which characterizes the weakest precondition for the postcondition $B$.
Let $S=(s_{1},\ldots,s_{m})$ be an $m$-tuple with each $s_{i}\subseteq[n]$, and consider a predicate $\mathscr{B}=(B_{s_{1}},\ldots,B_{s_{m}})$; we have
$M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}},$
where $0\leq B_{s_{i}}\leq I_{s_{i}}$.
Definition 4.5.
Given a fixed domain $S=(s_{1},\cdots,s_{m})$ and a quantum program $\mathbf{C}=\lambda\rho.U\rho U^{\dagger}$, operating on a density matrix, we say that $\mathscr{A}=(A_{s_{1}},\ldots,A_{s_{m}})$ is a local precondition of $\mathscr{B}=(B_{s_{1}},\ldots,B_{s_{m}})$ if
(7) ${\mathrm{Tr}}\left(M_{\mathscr{A}}\rho\right)\leq{\mathrm{Tr}}\left(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)\right)$
for input state $\rho$, where $M_{\mathscr{A}}=\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ and $M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}$, which would correspond to the (candidate) judgment
$\{M_{\mathscr{A}}\}\ C\ \{M_{\mathscr{B}}\}.$
We could use a QHL-like proof rule for a unitary, namely
[…]
To see that $M_{\mathscr{A}}\leq U^{\dagger}M_{\mathscr{B}}U$ leads to a valid QHL judgement—without requiring that the observables are bounded above by the identity—we observe that
(9) ${\mathrm{Tr}}\left(M_{\mathscr{A}}\rho\right)\leq{\mathrm{Tr}}\left(U^{\dagger}M_{\mathscr{B}}U\rho\right)={\mathrm{Tr}}\left(M_{\mathscr{B}}U\rho U^{\dagger}\right)={\mathrm{Tr}}\left(M_{\mathscr{B}}\,\llbracket\mathbf{C}\rrbracket(\rho)\right),$
where the inequality follows from 2.4 applied to the premise of Rule 8, and the first equality uses 2.2. Equation 9 gives us Equation 7, and thus 4.5 tells us that the conclusion of Rule 8 holds.
This definition naturally gives rise to a correctness judgment of the form $\{\mathscr{A}\}\ \mathbf{C}\ \{\mathscr{B}\}$ for quantum circuits. In addition to enabling backward reasoning, Rule 8 with 4.5 also supports forward reasoning by changing the premise of Rule 8 to
$UM_{\mathscr{A}}U^{\dagger}\leq M_{\mathscr{B}}.$
Unfortunately, it is not possible, in general, to use either version of Rule 8 algorithmically to compute the weakest precondition and the strongest postcondition, respectively. The issue is one of expressibility: can the answer be decomposed according to the chosen scheme $S=(s_{1},\ldots,s_{m})$? In general, this structural constraint of our candidate logic presents an obstacle. For instance, even when $U_{F}$ is a two-qubit unitary, the transformed observable
$U_{F}^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U_{F}$
may not admit a decomposition of the form $\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ for any choice of local observables $A_{s_{i}}$. That is, the equality
(10) $M_{\mathscr{A}}=U_{F}^{\dagger}M_{\mathscr{B}}U_{F}$
may have no solution for local predicates $\mathscr{A}$, even if $\mathscr{B}$ is given.
To see this limitation concretely, consider the following simple example where the circuit input is fixed as |$\ket{0}\ket{0}$.
Example 4.1.
The program is illustrated as the following CNOT circuit
Figure 1. CNOT circuit $\mathbf{C}$
Let us reason about circuit $\mathbf{C}$ using Equation 10. We choose $(s_{1},s_{2})=(\{q_{1}\},\{q_{2}\})$ and $M_{\mathscr{B}}=|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$.
$U_{\mathbf{C}}^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U_{\mathbf{C}}=|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|.$
We now show that there is no $A_{1}\geq 0$ and $A_{2}\geq 0$ that satisfies
(11) $|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|=A_{1}\otimes I+I\otimes A_{2}.$
We see that the left-hand side is orthogonal to $|1\rangle\langle 1|\otimes|0\rangle\langle 0|$. Therefore, we know that
${\mathrm{Tr}}[(|1\rangle\langle 1|\otimes|0\rangle\langle 0|)(A_{1}\otimes I+I\otimes A_{2})]=0$
$\Longrightarrow$ ${\mathrm{Tr}}[(|1\rangle\langle 1|A_{1})\otimes|0\rangle\langle 0|]+{\mathrm{Tr}}[|1\rangle\langle 1|\otimes(|0\rangle\langle 0|A_{2})]=0$
[…]
$\Longrightarrow$ $A_{1}=\lambda_{1}|0\rangle\langle 0|\quad\textrm{{and}}\quad A_{2}=\lambda_{2}|1\rangle\langle 1|$
$\Longrightarrow$ $A_{1}\otimes I+I\otimes A_{2}=\lambda_{1}|0\rangle\langle 0|\otimes I+\lambda_{2}I\otimes|1\rangle\langle 1|\neq|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|.$
{mdframed}
Observation: This calculation demonstrates that, in local reasoning, it is not always possible to express the weakest precondition.
We now show that under the QHL-like proof rule 8, sometimes, cannot prove some simple assertion. For instance, for the input |$\ket{00}$, the CNOT gate has no effect; therefore, the output state is also |$\ket{00}$. Thus, the following assertion holds:
(12) $\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}.$
To see this, we compute
$\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}$
[…]
$\Longleftrightarrow$ $|0\rangle\langle 0|\otimes I+|00\rangle\langle 00|+|11\rangle\langle 11|\leq|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$
$\Longleftrightarrow$ $|00\rangle\langle 00|+|11\rangle\langle 11|\leq I\otimes|0\rangle\langle 0|$
However, the final inequality is not valid. Therefore, according to Rule 8, Assertion 12 is not provable
⊬$\not\vdash\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}.$
4.4. Quantitative Judgments and Validity
In addition to the expressivity issues discussed in § 4.3, the predicates used in that section are not even capable of expressing the singleton quantum state |$\ket{0}^{\otimes n}$. This situation indicates that there is a mismatch with our requirements: 4.5 requires a triple to be valid for any state that satisfies the precondition predicate, whereas quantum programs typically focus on the specific input state |$\ket{0}^{\otimes n}$. For both of these reasons, we seek a more-expressive logic that still allows quantitative local reasoning.
Rather than using the predicate in 4.4 alone, we use quantum abstract interpretation (YP21, ) as an aid.
Definition 4.6.
A judgment is a triple of the form
$\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}$
for program $\mathbf{C}$ general predicates $\mathscr{A}$, $\mathscr{B}$ and projective predicates $\mathcal{P}$, $\mathcal{Q}$.
Definition 4.7 (Validity).
A judgment $\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}$ as defined in 4.6 is true if we have $\forall\rho\vDash^{QAI}\mathcal{P}$,
(13) $\llbracket\mathbf{C}\rrbracket(\rho)$ $\vDash^{QAI}\mathcal{Q},$
(14) ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ $\leq{{\mathrm{Tr}}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),$
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are the matrix representations of $\mathscr{A}$ and $\mathscr{B}$, as defined in 4.2.
Consider the special case $\mathcal{P}=\mathcal{Q}=\mathcal{I}$ with $\mathcal{I}:=(I_{s_{1}},\cdots,I_{s_{m}})$, we will write
$\{\mathscr{A}|\mathcal{I}\}\mathbf{C}\{\mathscr{B}|\mathcal{I}\}=\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}$
in which case this definition simplifies to 4.5.
We have that ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ only depends on the reduced density matrices of $\rho$ with respect to the systems $\set{s_{i}}$, as we know from Equation 6,
${\mathrm{Tr}}(M_{\mathscr{A}}\rho)=\sum_{i=1}^{m}{\mathrm{Tr}}\left(A_{s_{i}}\rho_{s_{i}}\right).$
In other words, ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ only depends on the reduced density matrices of $\rho$ on the systems $\set{s_{i}}$. Intuitively, this judgment tracks a linear function of the tuple of reduced density matrices to enable quantitative reasoning.
Moreover, $\rho_{s_{i}}\vDash P_{s_{i}}$ implies that $\rho_{s_{i}}=P_{s_{i}}\rho_{s_{i}}P_{s_{i}}$. This property implies
$\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}\rho_{s_{i}})=\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}P_{s_{i}}\rho_{s_{i}}P_{s_{i}})=\sum_{i=1}^{m}{\mathrm{Tr}}(P_{s_{i}}A_{s_{i}}P_{s_{i}}\rho_{s_{i}}).$
where the last step is due to 2.2.
Let $\mathscr{A}^{\prime}=(P_{s_{1}}A_{s_{1}}P_{s_{1}},\cdots,P_{s_{m}}A_{s_{m}}P_{s_{m}})$; we have
${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={{\mathrm{Tr}}}(M_{\mathscr{A}^{\prime}}\rho).$
That is,
$\{\mathscr{A}^{\prime}|\mathcal{P}\}\mathbf{skip}\{\mathscr{A}|\mathcal{P}\}\ \ \&\ \ \{\mathscr{A}|\mathcal{P}\}\mathbf{skip}\{\mathscr{A}^{\prime}|\mathcal{P}\}$
In other words, the qualitative insights from quantum abstract interpretation enhance the precision of the reasoning by providing “stronger” postconditions or “weaker” preconditions.555 Although it is not clear the order relationship between $\mathscr{A}$ and $\mathscr{A}^{\prime}$. At least we have ${\mathrm{Tr}}(A_{s_{i}})\geq{\mathrm{Tr}}((P_{s_{i}}P_{s_{i}})A_{s_{i}})\geq{\mathrm{Tr}}(P_{s_{i}}A_{s_{i}}P_{s_{i}})$, where we used Lemma 2.2.
4.5. Reduction
We now consider the relationship between the correctness $\models^{\rm QAI}$ in QAI and $\models$ in 4.7.
Theorem 4.1 (Reduction Principle).
Let the quantum program be a quantum circuit, and consider its behavior with respect to any input state. For any projective predicates $\mathcal{P}=(P_{s_{1}},\cdots,P_{s_{m}})$ and $\mathcal{Q}=(Q_{s_{1}},\cdots,Q_{s_{m}})$, $\mathcal{P}$ and $\mathcal{Q}$ can be regarded as the observables $\mathscr{P}:=(P_{s_{1}},\cdots,P_{s_{m}})$ and $\mathscr{Q}=(Q_{s_{1}},\cdots,Q_{s_{m}})$, respectively. Then we have the property
If $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ in the sense of 4.7, then $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$.
Proof.
Assume that $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ as defined in 4.7. Then for any $\rho$, we have
${\mathrm{Tr}}(M_{\mathscr{P}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$
Let us choose $\rho\models\mathcal{P}$. According to Lemma 2.5, we have
${\mathrm{Tr}}(M_{\mathscr{P}}\rho)={\mathrm{Tr}}(\sum_{i}P_{s_{i}}\rho_{s_{i}})=\sum_{i}{\mathrm{Tr}}(P_{s_{i}}\rho_{s_{i}})=\sum_{i}1=m.$
On the other hand,
${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$
$=$ ${\mathrm{Tr}}\left(\left(\sum_{i}Q_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\llbracket\mathbf{C}\rrbracket(\rho)\right)$
$=$ $\sum_{i}{\mathrm{Tr}}(Q_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))$
$\leq$ $\sum_{i}{\mathrm{Tr}}(I_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))$
$=$ $\sum_{i}1=m.$
where the inequality follows from 2.4 and $Q_{s_{i}}\leq I_{s_{i}}$.
Therefore, ${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$ and ${\mathrm{Tr}}(Q_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))=1$ for all $i$. According to Lemma 2.5, we know ${\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho)\vDash Q_{s_{i}}$, which implies
$\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\},$
where Equation 3 $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$ iff for all $\rho$, $\rho\vDash\mathcal{P}$ implies $\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q}$. ∎
4.6. Logical System with Soundness
The inference rules for program constructs in SAQR-QC are given in Figure 2.
$\begin{split}&\textsc{Skip}\quad\ \frac{}{\{\mathscr{A}|\mathcal{P}\}\mathbf{Skip}\{\mathscr{A}|\mathcal{P}\}}\\[2.84544pt] &\textsc{Unit}\quad\ \frac{\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})}{\{\mathscr{A}|\mathcal{P}\}\bar{q}:=U_{F}\left[\bar{q}\right]\{\mathscr{B}|U_{F}^{\sharp}(\mathcal{P})\}}\ \ \ \ \ \ \\[2.84544pt] &\textsc{Seq}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1}\{\mathscr{D}|\mathcal{R}\}\ \ \ \ \ \ \{\mathscr{D}|\mathcal{R}\}\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1};\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}\ \ \ \ \ \ \\[2.84544pt] &\textsc{Con}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\},\ \ \mathscr{D}\sqsubseteq\mathscr{A},\ \mathscr{B}\sqsubseteq\mathscr{E},\ \mathcal{R}\sqsubseteq\mathcal{P},\ \mathcal{Q}\sqsubseteq\mathcal{T}}{\{\mathscr{D}|\mathcal{R}\}\mathbf{C}\{\mathscr{E}|\mathcal{T}\}}\end{split}$
Figure 2. Inference rules for program constructs in SAQR-QC. We can use the proof rules for both forward reasoning or backward reasoning.
Theorem 4.2 (Soundness).
The proof system in Figure 2 is sound. That is, for quantum program $\mathbf{C}$, $\vdash\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}\ {\rm implies}\ \models\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}.$
We defer the proof to the appendix.
4.7. How to use the Rules for Scalable Reasoning
Among the proof rules in Figure 2, only the rule Unit presents a potential challenge to scalable reasoning. Specifically, the condition
(15) $\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\gamma(\mathcal{P})$
requires computing the projector $\gamma(\mathcal{P})$, which is generally intractable for systems with a large number of qubits. That is, given a postcondition $\mathscr{B}$ along with projective predicates $\mathcal{P}$—it remains unclear how to compute a suitable precondition $\mathscr{A}$ in a scalable manner.
To address this issue, we must allow for controlled approximations that trade some precision for tractability. In this subsection, we propose methods for applying the rule Unit while preserving scalability. The core idea is to construct inequalities over high-dimensional systems by composing inequalities over smaller, lower-dimensional subsystems. This compositional approach ensures that reasoning remains feasible, even as the number of qubits in the quantum system increases.
Warm-up exercise: Ignore QAI: As a warm-up exercise—which is potentially subject to the same kind of imprecision that we saw in § 4.3), but sometimes gives a precise-enough result (as we will see in § 5)—we can use the following inequality, in which we ignore all information from the QAI component (i.e., $\gamma(\mathcal{P})$):
(16) $M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\Longleftrightarrow\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}\leq U^{{\dagger}}\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}U.$
This inequality is linear, so it can be satisfied by fulfilling a set of inequalities in some number of smaller systems (indexed by $j$). We partition the domain of predicates $S=(s_{1},\ldots,s_{m})$, where each $s_{i}\subseteq[n]$. For $\cup_{j}T_{j}=\{1,2,\cdots,m\}$ with $T_{i}\cap T_{k}=\emptyset$ for all $i\neq k$, we can deal with $j$ inequalities as follows:
(17) $\textrm{for all $j$}\qquad\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\leq U^{{\dagger}}\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}U\Longrightarrow M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}}).$
Alternatively, for forward reasoning, Equation 17 takes the form
(18) $\textrm{for all $j$}\qquad U\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}U^{{\dagger}}\leq\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\Longrightarrow M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}}).$
A more precise approach: Ignoring $\gamma(\mathcal{P})$ away may lose too much precision. We can refine the approach of the warm-up exercise as follows. We again consider some number of smaller systems (indexed by $j$) by partitioning the domain of predicates: for $\cup_{j}T_{j}=\{1,2,\cdots,m\}$ with $T_{i}\cap T_{k}=\emptyset$ for all $i\neq k$. Clearly, we have
(19) $\gamma(\mathcal{P})=\cap_{i}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}\subseteq P_{j}::=\cap_{i\in T_{j}}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Now we can use the following inequalities in our reasoning steps:
(20) for all $\textrm{for all $j$}\qquad P_{j}\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)P_{j}\leq P_{j}U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)UP_{j}.$
Thus, our two basic reasoning tools will be Equation 20 along with Equations 17 and 18.
Equation 20 implies Equation 15. The proof is given in Appendix C.
5. Quantitative Reasoning of general GHZ circuit
This section analyzes a generalized GHZ circuit in which half of the gates are arbitrary single-qubit unitaries, resulting in a highly entangled and densely parameterized system that lies beyond the reach of the Gottesman–Knill theorem (Aaronson_2004, ). We evaluate several quantitative properties of the circuit’s output state using the techniques developed in this paper. In particular, we apply the “warm-up exercise” method from § 4.7 (Equation 17) to support scalable reasoning throughout the analysis. In the final step, we invoke Quantum Abstract Interpretation (QAI)—in particular, make use of 3.2—to infer that the output state resides within a two-dimensional subspace. We then apply our quantitative-reasoning framework to approximate the amplitudes of the basis vectors spanning this subspace. Notably, the resulting characterization of the output state is accurate up to a global phase factor when compared to the exact output vector. This example illustrates the point made near the end of § 1 that using SAQR-QC is not “cookbook.” The reason we obtain such a precise characterization of the output state is because of a careful choice of predicates.
Figure 3. GHZ circuit with one-qubit unitaries $U_{i}$.
We select the following domain ($(\{1,2\},\{2,3\},\cdots,\{n-1,n\})$.
STEP 1. We first select the precondition to be $\{\mathscr{A}|\mathcal{P}\}$ where
$\mathscr{A}$ $=(A_{1,2},A_{2,3},\cdots,A_{n-1,n})=(|++\rangle\langle++|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|)$
$\mathcal{P}$ $=(P_{1,2},P_{2,3},\cdots,P_{n-1,n})=(|00\rangle\langle 00|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|).$
One can verify that the initial state $|0\cdots 0\rangle\langle 0\cdots 0|\vDash\mathcal{P}$. We now use our proof rules to compute a postcondition for the GHZ circuit, given the precondition $\{\mathscr{A}|\mathcal{P}\}$.
After the first $H$ gate, we can use inequality 16 to derive the following:
$\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}\leq H^{{\dagger}}\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}H$
$\Longleftrightarrow$ $H\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}H^{{\dagger}}\leq\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}$
$\Longleftrightarrow$ $\sum_{i}H{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}H^{{\dagger}}\leq\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}$
$\Longleftrightarrow$ $\sum_{i}H{A_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}H^{{\dagger}}\leq\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}$
$\Longleftrightarrow$ $H{A_{1,2}}H^{{\dagger}}\otimes I_{[n]\setminus\{1,2\}}+\sum_{i>1}{A_{i,i+1}}\otimes HI_{[n]\setminus\{i,i+1\}}H^{{\dagger}}\leq\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}$
$\Longleftrightarrow$ $H{A_{1,2}}H^{{\dagger}}\otimes I_{[n]\setminus\{1,2\}}+\sum_{i>1}{A_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}\leq\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}.$
The final inequality above matches inequality 18 on a term-by-term basis. We can satisfy inequality 18 when each of the following single-term inequalities hold:
$H{A_{1,2}}H^{{\dagger}}\otimes I_{[n]\setminus\{1,2\}}\leq B_{1,2}\otimes I_{[n]\setminus\{1,2\}}\qquad\quad A_{i,i+1}\otimes I_{[n]\setminus\{i,i+1\}}\leq B_{i,i+1}\otimes I_{[n]\setminus\{i,i+1\}}\ \mathrm{for}\ i>1.$
In this case, each of the inequalities can be satisfied as an equality, yielding the postcondition $\{\mathscr{B}|\mathcal{Q}\}$
$\mathscr{B}=(B_{1,2},B_{2,3},\cdots,B_{n-1,n})=(HA_{1,2}H^{{\dagger}},A_{2,3},\cdots,A_{n-1,n})=(|0+\rangle\langle 0+|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|)$
$\mathcal{Q}=(Q_{1,2},Q_{2,3},\cdots,Q_{n-1,n})=(HP_{1,2}H^{{\dagger}},P_{2,3},\cdots,P_{n-1,n})=(|+0\rangle\langle+0|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|)$
After the first CNOT gate has been applied to qubits $q_{1}q_{2}$, we use the Unit Rule to obtain a postcondition $\{\mathscr{C}|\mathcal{R}\}:=\{(C_{1,2},C_{2,3},C_{3,4}\cdots,C_{n-1,n})|(R_{1,2},R_{2,3},R_{3,4}\cdots,R_{n-1,n})\}$. We use QAI to compute $R_{1,2}=|00\rangle\langle 00|+|11\rangle\langle 11|$, $R_{2,3}=|00\rangle\langle 00|+|10\rangle\langle 10|$, and $R_{i,i+1}=|00\rangle\langle 00|$ for $i>2$.
What is left to determine are suitable values for $C_{1,2}$ and $C_{2,3}$. According to the Unit Rule, $\mathscr{C}=(C_{1,2},C_{2,3},C_{3,4}\cdots,C_{n-1,n})$ satisfies
$\text{CNOT}_{1,2}(\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
$\Longleftrightarrow$ $\sum_{i}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
$\Longleftrightarrow$ $\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}+\sum_{i>2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
$\Longleftrightarrow$ $\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}+\sum_{i>2}(\text{CNOT}_{1,2}I_{[n]\setminus\{i,i+1\}}\text{CNOT}_{1,2}^{{\dagger}}\otimes B_{i,i+1})\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
$\Longleftrightarrow$ $\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}+\sum_{i>2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
At this point, we can satisfy inequality 18 by satisfying the following smaller inequalities:
$\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i\leq 2,\ \ }I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1},$
$\mathrm{and}\ \ (I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\leq(I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1})\ \mathrm{for}\ i>2.$
[…]
$\text{CNOT}_{1,2}(B_{1,2}\otimes I_{3}+I_{1}\otimes B_{2,3})\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $\text{CNOT}_{1,2}B_{1,2}\text{CNOT}_{1,2}^{{\dagger}}\otimes I_{3}+\text{CNOT}_{1,2}I_{1}\otimes B_{2,3}\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $\text{CNOT}_{1,2}|0+\rangle\langle 0+|\text{CNOT}_{1,2}^{{\dagger}}\otimes I_{3}+\text{CNOT}_{1,2}I_{1}\otimes|++\rangle\langle++|\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+\text{CNOT}_{1,2}(|0\rangle\langle 0|+|1\rangle\langle 1|)\otimes|++\rangle\langle++|\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+|0\rangle\langle 0|\otimes|++\rangle\langle++|+\text{CNOT}_{1,2}|1\rangle\langle 1|\otimes|++\rangle\langle++|\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+|0\rangle\langle 0|\otimes|++\rangle\langle++|+\text{CNOT}_{1,2}|1\rangle\langle 1|\otimes|+\rangle\langle+|\text{CNOT}_{1,2}^{{\dagger}}\otimes|+\rangle\langle+|\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+I_{1}\otimes|++\rangle\langle++|\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
where, in the third-to-last and last steps, we use the following facts:
[…]
Therefore, we find that the post-state $\mathscr{C}$ predicate is
(22) $(C_{1,2},C_{2,3},C_{3,4}\cdots,C_{n-1,n})=(|0+\rangle\langle 0+|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|).$
(which happens to be the same as the pre-state $\mathscr{B}$ predicate $(B_{1,2},B_{2,3},\cdots,B_{n-1,n})$), and the post-state $\mathcal{R}$ predicate is
$\mathcal{R}=(R_{1,2},R_{2,3},\cdots,R_{n-1,n})=(|00\rangle\langle 00|+|11\rangle\langle 11|,|00\rangle\langle 00|+|10\rangle\langle 10|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|).$
Equation 22 illustrates an advantage of our choice of predicates. Because of properties such as those given in Equation 21, Equation 16 remains invariant under the application of CNOT gates. This invariance allows us to derive the strongest postcondition, while preserving the local structure of the matrix representation of predicates. As a result, we were able to make choices that made the inequalities that we worked with tight (or saturated, i.e., satified as equalities), making it easier to determine the postcondition.
The right-hand side of Equation 22 continues to serve as the predicate of local observables, as reasoning continues about the remaining CNOT gates.
After applying the CNOT gate on $q_{1},q_{r}$ for some $r$, we can choose the postcondition to be $(\mathscr{D}|\mathcal{S})$ where
[…]
$\mathcal{S}$ $=(|00\rangle\langle 00|+|11\rangle\langle 11|,\cdots,|00\rangle\langle 00|+|11\rangle\langle 11|,|00\rangle\langle 00|+|10\rangle\langle 10|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|).$
At the right end of the circuit, after the application of $U_{1}\otimes U_{2}\otimes\cdots\otimes U_{n}$—where each $U_{i}$ is a single-qubit unitary—the locality structure of the predicates remains unchanged. Therefore, we can choose the postcondition to be $(\mathscr{F}\mid\mathcal{T})$, where
$\mathscr{F}$ $=(\beta_{1}\otimes\beta_{2},\cdots,\beta_{n-1}\otimes\beta_{n})$
$\mathcal{T}$ $=(\psi_{1}\otimes\psi_{2}+\phi_{1}\otimes\phi_{2},\cdots,\psi_{n-1}\otimes\psi_{n}+\phi_{n-1}\otimes\phi_{n})$
with $\beta_{i}=|\beta_{i}\rangle\langle\beta_{i}|,\psi_{i}=|\psi_{i}\rangle\langle\psi_{i}|,\ \ \phi_{i}=|\phi_{i}\rangle\langle\phi_{i}|$ and
|$\ket{\beta_{1}}$ $=U_{1}\ket{0},\ \ \ket{\beta_{i}}=U_{i}\ket{+}\ \ \forall i>1$
|$\ket{\psi_{i}}$ $=U_{i}\ket{0},\ \ \ \ket{\phi_{i}}=U_{i}\ket{1}\ \ \forall i\geq 1.$
Let us denote the output state as $\rho=|\Psi\rangle\langle\Psi|$. According to 4.7, our proof of $\mathscr{F}$ implies that
$\sum_{i=1}^{n-1}{\mathrm{Tr}}(|00\rangle\langle 00||++\rangle\langle++|)\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\beta_{i}\otimes\beta_{i+1})].$
That is
(23) $\frac{n-1}{4}\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\beta_{i}\otimes\beta_{i+1})].$
STEP 2. This step is similar to Step 1, except that we start with the precondition
$((|--\rangle\langle--|,|--\rangle\langle--|,\cdots,|--\rangle\langle--|)|\mathcal{P})$. We can compute a postcondition of the whole circuit as
$((\delta_{1}\otimes\delta_{2}\cdots,{\delta_{n-1}\otimes\delta_{n}})|\mathcal{T}),$
where $\mathcal{T}$ is the same as in Step 2, and $\delta_{i}=|\delta_{i}\rangle\langle\delta_{i}|,\ \ \ket{\delta_{1}}=U_{1}\ket{1},\ \ \ket{\delta_{i}}=U_{i}\ket{-}\ \ \forall i>1$. These conditions imply that
(24) $\frac{n-1}{4}\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\delta_{i}\otimes\delta_{i+1})].$
STEP 3 (QAI influence on QHL). According to Theorem 3.2, $\mathcal{Q}$ implies that the output state of the GHZ circuit is of form
|$\ket{\Psi}=a\ket{\psi_{1}\cdots\psi_{n}}+b\ket{\phi_{1}\cdots\phi_{n}}$
for complex numbers $|a|^{2}+|b|^{2}=1$.
According to $\langle\psi_{i}|\phi_{i}\rangle=0$, we have $\rho_{i,i+1}=|a|^{2}\psi_{i}\otimes\psi_{i+1}+|b|^{2}\phi_{i}\otimes\phi_{i+1}.$ Then
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\beta_{1}\otimes\beta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall 1<i$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\beta_{1}\otimes\beta_{2})]=0,\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall 1<i$
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\delta_{1}\otimes\delta_{2})]=0,\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall 1<i$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\delta_{1}\otimes\delta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall 1<i.$
Equations 23 and 24 imply
$\frac{n-1}{4}\leq|a|^{2}/2+\sum_{i=2}^{n-1}\frac{|a|^{2}+|b|^{2}}{4}=|a|^{2}/2+\frac{n-2}{4}\Longrightarrow\frac{1}{2}\leq|a|^{2}$
$\frac{n-1}{4}\leq|b|^{2}/2+\sum_{i=2}^{n-1}\frac{|a|^{2}+|b|^{2}}{4}=|b|^{2}/2++\frac{n-2}{4}\Longrightarrow\frac{1}{2}\leq|b|^{2}$
Together with $|a|^{2}+|b|^{2}=1$, we have $|a|^{2}=|b|^{2}=\frac{1}{2}$. In other words, there exists $\theta$ such that
(25) |$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{\psi_{1}\cdots\psi_{n}}+e^{i\theta}\ket{\phi_{1}\cdots\phi_{n}}).$
Equation 25 represents a closed-form expression for the circuit’s output with a single unknown real parameter. Such quantitative reasoning cannot be achieved with qualitative methods like QAI alone.
Moreover, the reasoning process scales with the number of qubits in the circuit: during the forward-reasoning process described above, for the reasoning steps carried out for each gate, the total size of the matrices that represent local observables and local projections in the pre- and post-conditions is always linear in the number of qubits. There are $O(n)$ gates; hence, the total amount of space needed to write down the SAQR-QC proof is $O(n^{2})$.
6. Quantum Phase Estimation
This section applies our proof system to analyze Quantum Phase Estimation (QPE), a foundational algorithm in quantum computing. QPE plays a pivotal role in applications such as Shor’s algorithm, quantum simulation, and other tasks requiring precise eigenvalue estimation—serving as a key source of quantum speedup over classical approaches.
We begin by verifying the correctness of the Quantum Fourier Transform (QFT) using Quantum Abstract Interpretation (QAI). Building upon the QAI results, we then apply our logic framework to reason about the Quantum Phase Estimation (QPE) circuit. In particular, we use our proof system to derive a lower bound on the success probability of QPE, demonstrating the power of our scalable quantitative reasoning approach.
To the best of our knowledge, QFT and QPE have not previously been analyzed using any scalable formal-reasoning framework. Our method provides a novel and tractable approach to reasoning about the correctness and quantitative behavior of QPE circuits.
6.1. Quantum Fourier transform
The QFT is the quantum analog of the discrete Fourier transform, central to algorithms like Shor’s for factoring, and quantum phase estimation. We first present a lossless local reasoning method for the quantum Fourier transform (QFT) based on Quantum Abstract Interpretation (QAI). Remarkably, for inputs in the computational basis, the concretized abstract output state produced by our analysis exactly matches the true output state. To the best of our knowledge, this is the first approach that demonstrates—within polynomial time—the ability to recover such precise information about the QFT, and, by extension, quantum phase estimation.
Figure 4. Quantum Fourier transform. Swap gates at the end of the circuit that reverse the order of the qubits are not shown.
The quantum gates used in the circuit are the Hadamard gate and the phase gate $R_{m}$, $R_{m}$ does not belong to the Clifford group for $m>2$.
${H={\frac{1}{\sqrt{2}}}{\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}}\qquad{\text{and}}\qquad R_{m}={\begin{pmatrix}1&0\\ 0&e^{2\pi i/2^{m}}\end{pmatrix}}}$
We use the $\psi=|\psi\rangle\langle\psi|$ for pure state |$\ket{\psi}$,$0.x_{1}x_{2}\cdots x_{n}=\sum_{i=1}^{n}\frac{x_{i}}{2^{i}}$, and the following notation
(26) |$\ket{\psi_{x}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{2\pi i0.x}\ket{1}).$
[…]
$\mathcal{P}=(P_{1},P_{2},\cdots,P_{n})=(|j_{1}\rangle\langle j_{1}|,|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|).$
In Appendix C, we prove the following using QAI
(27) $\vDash^{QAI}\{\mathcal{P}\}QFT\{(\psi_{j_{n}},\psi_{j_{n-1}j_{n}},\cdots,\psi_{j_{2}\cdots j_{n}},\psi_{j_{1}\cdots j_{n}})\}$
The postcondition derived from QAI is an abstract state represented as a tuple of density matrices corresponding to pure quantum states. By applying the concretization function from 3.2, we infer that the concrete state lies in the subspace
$\psi_{j_{n}}\otimes\psi_{j_{n-1}j_{n}}\otimes\cdots\otimes\psi_{j_{2}\cdots j_{n}}\otimes\psi_{j_{1}\cdots j_{n}}.$
Because the space so defined is a 1-dimensional subspace, it follows that the density matrix of the output state must exactly equal the pure-state projection onto this vector—that is, $\psi_{j_{n}}\otimes\psi_{j_{n-1}j_{n}}\otimes\cdots\otimes\psi_{j_{2}\cdots j_{n}}\otimes\psi_{j_{1}\cdots j_{n}}$.
During the forward-reasoning process described above, for the reasoning steps carried out for each gate, the total size of the matrices that represent local observables and local projections in the pre- and post-conditions is always linear in the number of qubits. There are $O(n^{2})$ gates; hence, the total amount of space needed to write down the proof is $O(n^{3})$.
6.2. Quantum Phase Estimation
In this section, we present a quantitative local analysis of the Quantum Phase Estimation (QPE) algorithm by combining both backward and forward reasoning techniques. We begin by decomposing the QPE algorithm into three constituent circuits and apply SAQR-QC to reason about each component individually. For an unknown phase $\theta$ and a given constant $k$, SAQR-QC can formally establish that—with probability at least $\frac{4}{\pi^{2}}$—the QPE algorithm produces an output bitstring whose last $k$ bits constitute the optimal $k$-bit approximation to the least significant $k$ bits of any $n$-bit binary representation/approximation of $\theta$.
Figure 5. Quantum Phase Estimation: $U\ket{\psi}=e^{i\theta}\ket{\psi}$. We only consider the circuit without the measurements.
We denote by $\mathbf{C_{1}}$ the segment of the program that precedes the application of the inverse Quantum Fourier Transform, $\mathcal{QFT}^{{\dagger}}$. The $\mathcal{QFT}^{{\dagger}}$ operation can be decomposed into two parts: the initial sequence of swap gates that reverses the order of the qubits, and the subsequent controlled rotation gates that implement the core of the inverse Fourier transform. We use $\mathbf{C_{2}}$ to refer specifically to the subcircuit following the swap gates within $\mathcal{QFT}^{{\dagger}}$.
Our analysis focuses on the least significant $k$ qubits, and proceeds in three steps. We will rely on the notation introduced in Equation 36, and apply the QAI-based reasoning framework introduced in § 4.7 to verify the behavior of $\mathbf{C_{1}}$.
Step 1: Reasoning about $\mathbf{C_{2}}$. Let $U$ be the circuit corresponding to $\mathbf{C_{2}}$. For $\mathbf{C_{2}}$, we choose the domain of the predicate as the first $k$ qubits of the first $n$ qubits, together with the last $m$ qubits—the qubits that $U$ acts upon. We use backward reasoning to show
(28) $\{\mathscr{A}|\mathcal{I}\}\mathbf{C_{2}}\{(|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)|\mathcal{I}\}.$
with $\mathscr{A}=(|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|)$,and |$\ket{\tau}=\ket{\psi_{j_{n}}}\otimes\cdots\otimes\ket{\psi_{j_{n-k+1}\cdots j_{n}}}$. Then,
$M_{\mathscr{A}}=I_{1,\cdots,n-k}\otimes|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|.$
Equation 28 is equivalent to
$M_{\mathscr{A}}\leq$ $U^{{\dagger}}[I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|\otimes|\psi\rangle\langle\psi|]U$
$=$ $U^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)U\otimes|\psi\rangle\langle\psi|.$
To see this, we observe
$H(1)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)H(1)^{{\dagger}}$
$=$ $CR_{n}(n,1)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)CR_{2}(n,1)^{{\dagger}}$
$=$ $CR_{2}(2,1)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)CR_{2}(2,1)^{{\dagger}}$
$=$ $H(n-k+2)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)H(n-k+2)^{{\dagger}}$
$=$ $I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|.$
Let $V=H(1)^{-1}CR_{2}(2,1)^{-1}CR_{3}(3,1)^{-1}\cdots H(n-k+2)^{-1}$ means
$V^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)V=I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|$
We let $U=VU_{k}$ with $U_{k}$ being the sub-circuit which only applied on the last $k$-qubits, i.e.,
$U_{k}=H(n-k+1)^{-1}\cdots H(n-1)^{-1}CR_{2}(n,n-1)^{-1}H(n)^{-1},$
Then
$U^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)U\otimes|\psi\rangle\langle\psi|$
$=$ $U_{k}^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)U_{k}\otimes|\psi\rangle\langle\psi|$
$=$ $I_{1,\cdots,n-k}\otimes U_{k}^{{\dagger}}|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|U_{k}\otimes|\psi\rangle\langle\psi|$
$U_{k}^{{\dagger}}$ is the standard Quantum Fourier transform applied on input $|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|$. Performing direct matrix multiplication (or according to the last subsection of reasoning about the Quantum Fourier transform), we know that
$U_{k}^{{\dagger}}|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|U_{k}=|\tau\rangle\langle\tau|$
This argument proves Equation 28.
Remark:
Matrix multiplication will be efficient for constant dimension. Here, the dimension of matrices is $2^{k}$, which is a constant independent of $n$.
Step 2: Reasoning about SWAP gates. Now the precondition of Equation 28, $\{\mathscr{A}|\mathcal{I}\}$, becomes the postconditon of the SWAP gates. Since the action of the SWAP gates only changes the last $k$ qubits into the first $k$ qubits, we may just write the precondition of the SWAP gates as $\{\mathscr{A}^{\prime}|\mathcal{I}\}$ with $\mathscr{A}^{\prime}=(|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|)$ on the first $k$ qubits and the last $m$ qubits. These arguments prove the following
(29) $\{\mathscr{A}^{\prime}|\mathcal{I}\}\mathbf{S}\{\mathscr{A}|\mathcal{I}\}$
where we use $\mathbf{S}$ to denote the SWAP gates.666 We change the abstract domain for simplicity of presentation. This change will not affect our statement’s correctness because the SWAP gates’ action is clear. If we want to fix the abstract domain, we can consider $(s_{1},\cdots,s_{m})$ with $s_{1}$ being the last $k$ qubits together with the last $m$ qubits; $s_{2}$ being the result of applying the first SWAP gate on $s_{1}$; $\cdots$; $s_{m}$ being the result of applying the last SWAP gate on $s_{m-1}$.
Step 3: Reasoning about $\mathbf{C_{1}}$. For $\mathbf{C_{1}}$, we choose the domain of the predicate as the first $k$ qubits of the first $n$ qubits together with the last $m$ qubits—the qubits on which $U$ applied. We will use our proof rules to show
(30) $\{r|0\cdots 0\rangle\langle 0\cdots 0|\otimes|\psi\rangle\langle\psi|)|(|0\cdots 0\rangle\langle 0\cdots 0|\otimes|\psi\rangle\langle\psi|\}\mathbf{C_{1}}\{\mathscr{A}^{\prime}|\mathcal{P}\}$
where
$r$ $=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]$
$\mathcal{P}$ $=(|\omega\rangle\langle\omega|\otimes|\psi\rangle\langle\psi|)$
|$\ket{\omega}$ $=\frac{1}{2^{k/2}}(\ket{0}+e^{2\pi i2^{n-1}\theta}\ket{1})\otimes\cdots\otimes(\ket{0}+e^{2\pi i2^{n-k}\theta}\ket{1}).$
We first use quantum abstract interpretation (YP21, ) to prove
(31) $\{\cdot|(|0\cdots 0\rangle\langle 0\cdots 0|\otimes|\psi\rangle\langle\psi|)\}\mathbf{C_{1}}\{\cdot|\mathcal{P}\}.$
After applying the first $H$ gates, the post-condition becomes
$\{\cdot|(|+\cdots+\rangle\langle+\cdots+|\otimes|\psi\rangle\langle\psi|\}.$
For $CU^{2^{n-1}}\cdots CU^{2^{n-k}}$, direct matrix computation leads to the post-condition $\{\cdot|\mathcal{P}\}$.
For $CU^{2^{n-k+1}}\cdots CU^{2^{0}}$, there will be no change to the observable on the first $k$ qubits. For example, for the first gate $CU^{2^{0}}$, which is $CU$, we have the post-condition $\{\cdot||+\cdots+\rangle\langle+\cdots+|\otimes|\psi\rangle\langle\psi|\}$ by observing
$\operatorname{supp}[{\mathrm{Tr}}_{n}CU(|\omega\rangle\langle\omega|\otimes I_{n}\otimes|\psi\rangle\langle\psi|)CU^{{\dagger}}]$
$=$ $\operatorname{supp}[{\mathrm{Tr}}_{n}CU(|\omega\rangle\langle\omega|\otimes(|0\rangle\langle 0|+|1\rangle\langle 1|)\otimes|\psi\rangle\langle\psi|)CU^{{\dagger}}]$
$=$ $\operatorname{supp}[2|\omega\rangle\langle\omega|\otimes|\psi\rangle\langle\psi|]$
$=$ $|\omega\rangle\langle\omega|\otimes|\psi\rangle\langle\psi|.$
Similarly for $CU^{2^{n-k+1}}\cdots CU^{2^{0}}$. This argument proves Equation 31.
To prove Equation 30, we observe
$|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|(r|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|)|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|$
$=$ $r|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|;$
and
$|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|[U(C_{1})^{{\dagger}}(|\tau\rangle\langle\tau|\otimes I\otimes|\psi\rangle\langle\psi|)U(C_{1})]|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|$
$=$ $|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|[V(C_{1})^{{\dagger}}(|\tau\rangle\langle\tau|\otimes I\otimes|\psi\rangle\langle\psi|)V(C_{1})]|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|$
$=$ $x|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|;$
where
$V(C_{1})$ $=CU^{2^{n-k}}\cdots CU^{2^{n-1}}$
$x$ $=\langle 0\cdots 0\psi|V(C_{1})^{{\dagger}}|\tau\psi\rangle\langle\tau\psi|\ket{0\cdots 0\psi}$
$={\mathrm{Tr}}[(|\tau\psi\rangle\langle\tau\psi|)V(C_{1})|0\cdots 0\psi\rangle\langle 0\cdots 0\psi|V(C_{1})^{{\dagger}}]$
$={\mathrm{Tr}}[|\tau\psi\rangle\langle\tau\psi||\omega\psi\rangle\langle\omega\psi|]$
$=|\langle\tau|\omega\rangle|^{2}$
$=\Pi_{t=1}^{k}|\frac{(\langle 0|+e^{-2\pi i0.j_{n-t+1}\cdots j_{n}}\langle 1|)(\ket{0}+e^{2\pi i2^{n-t}\theta}\ket{1}}{2}|^{2}$
$=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]$
$=r.$
$r$ $=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]=\Pi_{t=1}^{k}\frac{\sin^{2}[2(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]}{4\sin^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]}$
$=\Pi_{t=1}^{k}\frac{\sin^{2}[2^{n-t+1}\theta-0.j_{n-t+2}\cdots j_{n})\pi]}{4\sin^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]}=\frac{\sin^{2}(2^{n}\theta\pi)}{4^{k}\sin^{2}[(2^{n-k}\theta-0.j_{n-k+1}\cdots j_{n})\pi]}.$
Let $U\ket{\psi}=e^{i\theta}\ket{\psi}$, and $\theta=\frac{a}{2^{n}}+\epsilon$ with $-\frac{1}{2^{n+1}}\leq\epsilon\leq\frac{1}{2^{n+1}}$ and $a$ is an integer with binary representation $a_{1}a_{2}\cdots a_{n}$. For $j_{n-k}\cdots j_{n}=a_{n-k}\cdots a_{n}$, we have
$r$ $=\frac{\sin^{2}(2^{n}\theta\pi)}{4^{k}\sin^{2}[(2^{n-k}\theta-0.j_{n-k+1}\cdots j_{n})\pi]}=\frac{\sin^{2}(2^{n}\epsilon\pi)}{4^{k}\sin^{2}(2^{n-k}\epsilon\pi)}\geq\frac{|2\cdot 2^{n}\epsilon|^{2}}{4^{k}\cdot|2^{n-k}\epsilon\cdot\pi|^{2}}\geq\frac{4}{\pi^{2}}.$
Together with the Seq Rule and the Con Rule, the proved result can be interpreted as follows: for any constant $k$, the last $k$ bits of the output will—with probability at least $\frac{4}{\pi^{2}}$—match the best $k$-bit binary approximation to the least significant $k$ bits of the phase $\theta$, provided that the first $k$ input qubits are initialized to |$\ket{0,\ldots,0}$, regardless of the state of the remaining $n-k$ qubits.
7. Related Work
Quantum Hoare logic. Researchers have developed a variety of quantum Hoare logics. (RA17, ) compared three such logics, namely (CMS, ; 10.1007/978-3-642-10622-4_7, ; Ying11, ). We can divide quantum Hoare logic into expectation-based and satisfaction-based approaches. Following the approach in the seminal paper by D’Hondt and Panangaden (DP06, ), the expectation-based approaches in (Ying11, ; BHY19, ; li2019quantum, ; feng2021quantum, ) take positive operators as assertions for quantum states. This approach enables the expectation that a quantum state $\rho$ satisfies an assertion $M$ to be defined as ${\mathrm{Tr}}(M\rho)$. In contrast, the satisfaction-based logics (ZYY19, ; Unruh, ; Unruh2, ) regard subspaces of the Hilbert space as assertions. This approach enables the assertion that a quantum state $\rho$ satisfies an assertion $P$ to be defined as the support (the image space of linear operators) of $\rho$ is included in $P$. All the mentioned papers represent a predicate $M$ or $P$ as a $2^{n}\times 2^{n}$ dimensional matrix. In other words, they do not provide an efficiently computable quantum logic.
Quantum Separation logic. Quantum separation logic in (10.1109/LICS52264.2021.9470673, ) enables local reasoning for quantum computation using a quantum interpretation of Bunched Implications (o'hearn_pym_1999, ). The quantum separation logic in (10.1145/3498697, ) supports classical variables and quantum qubits’ dynamic allocation/deallocation. In these works, the separating conjunction is defined as a tensor product, i.e., quantum independence. This requirement significantly restricts the applicability of these logics.
Quantum abstract interpretation. (YP21, ) presented an approach to quantum abstract interpretation for reasoning about quantum circuits, using the satisfaction-based approach. Other works, such as (Bichsel_2023, ), investigate the abstract interpretation of quantum programming using variants of the Gottesman-Knill theorem.
Symbolic abstraction, strongest consequence, and weakest sufficient condition The inexpressibility issues discussed in § 4.3 are a manifestation of the constraints that one faces when working with an “impoverished” logic (or logic fragment). These issues have been studied in the context of abstract interpretation as what is (now) called the symbolic-abstraction problem (DBLP:conf/vmcai/RepsSY04, ; DBLP:conf/cav/ThakurR12, ), and phrased in purely logical terms as the strongest-consequence problem (DBLP:conf/vmcai/RepsT16, , §5), as follows: {mdframed}[innerleftmargin = 3pt, innerrightmargin = 3pt, skipbelow=-0.5em] Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the strongest formula $\psi\in\mathcal{L}^{\prime}$ such that $\varphi\vDash\psi$. The strongest-consequence problem naturally arises in approximate forwards reasoning, to over-approximate a postcondition. The discussion in § 4.3 concerned backwards reasoning for which one faces the dual problem, the weakest sufficient-condition problem: {mdframed}[innerleftmargin = 3pt, innerrightmargin = 3pt, skipbelow=-0.5em] Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the weakest formula $\chi\in\mathcal{L}^{\prime}$ such that $\chi\vDash\varphi$.
As we saw in § 4.3, the strongest consequence or weakest sufficient condition may not always be expressible in $\mathcal{L}^{\prime}$, in which case one has to fall back on finding what one hopes is a suitably strong consequence or a suitably weak sufficient condition, respectively. Scherpelz et al. (DBLP:conf/pldi/ScherpelzLC07, ) presented a best-effort method for computing sufficient conditions as part of an algorithm for creating abstract transformers for use with parameterized predicate abstraction (DBLP:conf/birthday/Cousot03, ). Their method performs weakest liberal precondtion (WLP) of a post-state predicate with respect to a concrete transformer $\tau$, and then uses heuristics to identify combinations of pre-state predicates that entail the WLP value.
Reasoning about Shor’s factoring.
The QFT circuit contains many non-Clifford gates, so variants of the Gottesman-Knill theorem do not directly apply. Shor’s algorithm has been formally verified in (feng2021quantum, ; https://doi.org/10.48550/arxiv.2204.07112, ). The former uses an expectation-based approach with classical variables, while the latter employs the Coq proof assistant and the Small Quantum Intermediate Representation (sQIR) (Hicks21, ). Since Coq operates symbolically, the reasoning about quantum phase estimation in (https://doi.org/10.48550/arxiv.2204.07112, ) is symbolic.
Symbolic verification. There has been work to extend symbolic-verification techniques to the quantum domain (10.1145/3725725, ; CACM:CCLLTY25, ), using logic- and automata-based techniques developed for symbolic verification of classical programs to analyze the correctness of quantum programs. In contrast, SAQR-QC is not symbolic in nature. Reasoning steps in SAQR-QC can involve matrix multiplications and other mathematical operations on specific values of specific sizes.
SAQR-QC also does not support parameterized reasoning about families of circuits parameterized on the number of qubits. On the contrary, SAQR-QC can be used to reason about a specific circuit with a specific number of qubits.
Packing of variables in abstract domains The motivation for using local observables and local projections defined with respect to a tuple $(s_{1},\cdots,s_{m})$ of sets of qubit indexes is to make reasoning scalable: each reasoning step involves only a small number of qubits. This idea is similar to the idea of “packing” variables in numeric domains, as used in Astrée (DBLP:conf/pldi/BlanchetCCFMMMR03, ): a program’s numeric variables are placed in sets (“packs”) so that each abstract transformer of a numeric abstract domain can operate on a single pack at a time. As with our qubit sets, each variable can be in multiple packs.
8. Conclusion and Future work
This paper introduces a quantitative local-reasoning framework for quantum circuits, bridging a gap in quantum program analysis. By tracking tuples of reduced-density matrices through linear functions, our method enables efficient verification of quantum circuits, including Quantum Phase Estimation, Quantum Fourier Transform, and GHZ circuits with non-Clifford gates. Unlike existing scalable frameworks limited to Clifford circuits (GottesmanKnill, ), ours extends to non-Clifford gates, significantly broadening applicability. Below, we outline future research directions for efficient reasoning about quantum programs.
One direction is to extend the logic to include multiple tuples of local observables. Specifically, we consider judgments of the form:
(32) $\{(\mathscr{A}_{1},\cdots,\mathscr{A}_{k})\,|\,\mathcal{P}\}\ \mathbf{C}\ \{(\mathscr{B}_{1},\cdots,\mathscr{B}_{k})\,|\,\mathcal{Q}\},$
where each $\mathscr{A}_{i}$ and $\mathscr{B}_{i}$ is a tuple of observables. In this setting, rather than a single inequality, a judgment yields a collection of inequalities, enabling more fine-grained and expressive reasoning about the program’s behavior. We have seen the power of this idea in the GHZ example in § 5.
It would be interesting to extend our methods to general quantum programs with classical control, such as those described in (Ying11, ; ZYY19, ). Recent papers (10.1145/3453483.3454029, ; Wu19, ; ZYY19, ) have demonstrated how to use logic to reason about noise in quantum programs. We leave the design of an efficient logic for noisy quantum programs to future work.
It is also highly compelling to explore the automation of our methods for reasoning about quantum programs. One major obstacle, shared with the automation of QAI, is the need for effective heuristics to automatically select appropriate qubit domains for analysis. An additional challenge arises from the need to resolve matrix inequalities involving positive semidefiniteness. While semidefinite constraints are common in optimization, our setting differs significantly: rather than optimizing a scalar objective function, we aim to synthesize a matrix—specifically, a postcondition observable or predicate—that satisfies a semidefinite inequality.
This requirement introduces structural difficulties. The set of positive semidefinite matrices is not a lattice under the Löwner partial order: given two such matrices $A$ and $B$, a least upper bound (i.e., a “maximum” matrix) may not exist within the set. Consequently, unlike in classical predicate synthesis or optimization, we cannot rely on lattice-theoretic fixed-point techniques to guide the construction of such matrices. Addressing this issue is essential for developing scalable and principled automated reasoning tools for quantum programs.
Appendix A Appendix A: Lemmas and Proofs of Lemmas
Lemma A.1.
$P\subseteq Q\Longrightarrow PQ=QP=P.$
Lemma A.2.
$A\leq B\Longrightarrow PAP\leq PBP.$
2.1. Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$,
${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right).$
Proof.
We express $\rho$ as
$\rho=\sum_{i,j}\rho_{i,j}\otimes\ket{i}\bra{j},$
where $\rho_{i,j}$ are operators on subsystem $s$, and {|$\{\ket{i}\}$ is an orthonormal basis of subsystem $[n]\setminus s$. Then:
${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)$ $={\mathrm{Tr}}\left(\sum_{i,j}(A_{s}\rho_{i,j})\otimes\ket{i}\bra{j}\right)$
$=\sum_{i}{\mathrm{Tr}}(A_{s}\rho_{i,i})$
$={\mathrm{Tr}}\left(A_{s}\sum_{i}\rho_{i,i}\right)$
$={\mathrm{Tr}}\left(A_{s}\rho_{s}\right),$
where $\rho_{s}={\mathrm{Tr}}_{[n]\setminus s}(\rho)=\sum_{i}\rho_{i,i}$.
Appendix B Appendix B: Proof of Theorem 4.2
Proof.
We prove the soundness of each rule.
Rule Skip
$\textsc{Skip}\quad\ \frac{}{\{\mathscr{A}|\mathcal{P}\}\mathbf{Skip}\{\mathscr{A}|\mathcal{P}\}}$
For any $\rho\vDash^{QAI}\mathcal{P}$, we have $\llbracket\mathbf{Skip}\rrbracket(\rho)=\rho\vDash^{QAI}\mathcal{P}$. Moreover,
${\mathrm{Tr}}[M_{\mathscr{A}}(\rho)]\leq{\mathrm{Tr}}[M_{\mathscr{A}}(\llbracket\mathbf{Skip}\rrbracket(\rho))].$
Rule Unit
$\textsc{Unit}\quad\ \frac{\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})}{\{\mathscr{A}|\mathcal{P}\}{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\{\mathscr{B}|U_{F}^{\sharp}(\mathcal{P})\}}$
For $\rho\vDash^{QAI}\mathcal{P}$, using Theorem 3.1, we have
$\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\rho)=U_{F}\rho U_{F}^{{\dagger}}\vDash^{QAI}U_{F}^{\sharp}(\mathcal{P}).$
Moreover, we have $\rho\vDash\gamma(\mathcal{P})$; that is, $\rho=\gamma(\mathcal{P})\rho\gamma(\mathcal{P})$. According to Lemma 2.2, we know
${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={{\mathrm{Tr}}}(M_{\mathscr{A}}\gamma(\mathcal{P})\rho\gamma(\mathcal{P}))={\mathrm{Tr}}(\rho\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P}))$
On the other hand, Lemma 2.2 also implies
${{\mathrm{Tr}}}(M_{\mathscr{B}}\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\rho))$
$=$ ${{\mathrm{Tr}}}[M_{\mathscr{B}}\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\gamma(\mathcal{P})\rho\gamma(\mathcal{P})]$
$=$ ${\mathrm{Tr}}[M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})\rho\gamma(\mathcal{P})U_{F}^{{\dagger}}]$
$=$ ${\mathrm{Tr}}[\rho\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})].$
Therefore, the condition
$\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\gamma(\mathcal{P})$
implies that, for $\rho\vDash^{QAI}\mathcal{P}$, we have
${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}(\rho\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P}))\leq{\mathrm{Tr}}[\rho\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})]\leq{{\mathrm{Tr}}}(M_{\mathscr{B}}\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\rho))$
by invoking Lemma 2.2.
This argument proves Rule Unit.
Rule Seq
$\textsc{Seq}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1}\{\mathscr{D}|\mathcal{R}\}\ \ \ \ \ \ \{\mathscr{D}|\mathcal{R}\}\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1};\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}$
For $\rho\vDash^{QAI}\mathcal{P}$, we have
$\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1}\{\mathscr{D}|\mathcal{R}\}\Longrightarrow\llbracket\mathbf{C}_{1}\rrbracket(\rho)\vDash^{QAI}\mathcal{R},\ \ \ {\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{D}}(\llbracket\mathbf{C}_{1}\rrbracket(\rho))].$
According to $\{\mathscr{D}|\mathcal{R}\}\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}$ and $\llbracket\mathbf{C}_{1}\rrbracket(\rho)\vDash^{QAI}\mathcal{R}$, we obtain that
$\llbracket\mathbf{C}_{2}\rrbracket(\llbracket\mathbf{C}_{1}\rrbracket(\rho))\vDash^{QAI}\mathcal{Q}$
${\mathrm{Tr}}[M_{\mathscr{D}}(\llbracket\mathbf{C}_{1}\rrbracket(\rho))]\leq{\mathrm{Tr}}[\llbracket M_{\mathscr{B}}(\mathbf{C}_{2}\rrbracket(\llbracket\mathbf{C}_{1}\rrbracket(\rho)))]$
According to the fact that $\llbracket\mathbf{C}_{1};\mathbf{C}_{2}\rrbracket(\rho))=\llbracket\mathbf{C}_{2}\rrbracket(\llbracket\mathbf{C}_{1}\rrbracket(\rho))$, we know that if $\rho\vDash^{QAI}\mathcal{P}$,
$\llbracket\mathbf{C}_{1};\mathbf{C}_{2}\rrbracket(\rho))\vDash^{QAI}\mathcal{Q}$
${\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{B}}(\llbracket\mathbf{C}_{1};\mathbf{C}_{2}\rrbracket(\rho))]$
This argument proves Rule Seq.
Rule Con
$\textsc{Con}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\},\ \ \mathscr{D}\sqsubseteq\mathscr{A},\ \mathscr{B}\sqsubseteq\mathscr{E},\ \mathcal{R}\sqsubseteq\mathcal{P},\ \mathcal{Q}\sqsubseteq\mathcal{T}}{\{\mathscr{D}|\mathcal{R}\}\mathbf{C}\{\mathscr{E}|\mathcal{T}\}}$
For any $\rho\vDash^{QAI}\mathcal{R}$, we have
$\rho\vDash^{QAI}\mathcal{R}\sqsubseteq\mathcal{P}.$
Moreover,
$\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}\Longrightarrow\llbracket\mathbf{C}\rrbracket(\rho)\vDash^{QAI}\mathcal{Q}\sqsubseteq\mathcal{T},\ \ {\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{B}}(\llbracket\mathbf{C}\rrbracket(\rho))].$
According to Lemma 4.1, we have
$M_{\mathscr{D}}\leq M_{\mathscr{A}},\ \ M_{\mathscr{B}}\leq M_{\mathscr{E}}$
$\Longrightarrow$ ${\mathrm{Tr}}[M_{\mathscr{D}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{B}}(\llbracket\mathbf{C}\rrbracket(\rho))]\leq{\mathrm{Tr}}[M_{\mathscr{E}}(\llbracket\mathbf{C}\rrbracket(\rho))].$
This argument proves the correctness of Rule Con. ∎
Appendix C Appendix C: Equation (20) implies Equation (15)
(33) for all $j$ $\gamma(\mathcal{P})P_{j}\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)P_{j}\gamma(\mathcal{P})$ $\leq\gamma(\mathcal{P})P_{j}U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)UP_{j}\gamma(\mathcal{P})$
(34) $\Longrightarrow$ for all $j$ $\gamma(\mathcal{P})\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\gamma(\mathcal{P})$ $\leq\gamma(\mathcal{P})U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\gamma(\mathcal{P})$
$\Longrightarrow$ $\sum_{j}\gamma(\mathcal{P})\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\gamma(\mathcal{P})$ $\leq\sum_{j}\gamma(\mathcal{P})U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\gamma(\mathcal{P})$
(35) $\Longrightarrow$ $\gamma(\mathcal{P})\left(\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\gamma(\mathcal{P})$ $\leq\gamma(\mathcal{P})U^{{\dagger}}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\gamma(\mathcal{P})$
where in going from Equation 20 to Equation 33, we used A.2; and in going from Equation 33 to Equation 34, we used Equation 19 and A.1.
Appendix D Appendix D: Reasoning about QFT using QAI
The quantum gates used in the circuit are the Hadamard gate and the phase gate $R_{m}$, $R_{m}$ does not belong to the Clifford group for $m>2$.
${H={\frac{1}{\sqrt{2}}}{\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}}\qquad{\text{and}}\qquad R_{m}={\begin{pmatrix}1&0\\ 0&e^{2\pi i/2^{m}}\end{pmatrix}}}$
We use the $\psi=|\psi\rangle\langle\psi|$ for pure state |$\ket{\psi}$,$0.x_{1}x_{2}\cdots x_{n}=\sum_{i=1}^{n}\frac{x_{i}}{2^{i}}$, and the following notation
(36) |$\ket{\psi_{x}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{2\pi i0.x}\ket{1}).$
We choose the domain ($(\{1\},\{2\},\cdots,\{n\})$, and the precondition to be
$\mathcal{P}=(P_{1},P_{2},\cdots,P_{n})=(|j_{1}\rangle\langle j_{1}|,|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|).$
At the beginning of the program, according to $\beta_{k}=|j_{k}\rangle\langle j_{k}|\models_{p}P_{k}$, we have the input
[…]
After applying the first $H$ gate on $q_{1}$, we compute the postcondition, which becomes
$(H|j_{1}\rangle\langle j_{1}|H,|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)=(\psi_{j_{1}},|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
where |$\ket{\psi_{j_{1}}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{2\pi i0.j_{1}}\ket{1})$ and $\psi_{j_{1}}=|\psi_{j_{1}}\rangle\langle\psi_{j_{1}}|$, by notifying $e^{2\pi i0.j_{1}}=-1$ if $j_{1}=0$, otherwise, $e^{2\pi i0.j_{1}}=1$.
Applying the controlled-$R_{2}$ gate, we compute
$P_{1,2}:=\psi_{j_{1}}\otimes I_{2}\cap I_{1}\otimes|j_{2}\rangle\langle j_{2}|=\psi_{j_{1}}\otimes|j_{2}\rangle\langle j_{2}|.$
The postcondition becomes
$(\operatorname{supp}({\mathrm{Tr}}_{2}CR_{2}A_{1,2}CR_{2}^{{\dagger}}),\operatorname{supp}({\mathrm{Tr}}_{1}CR_{2}A_{1,2}CR_{2}^{{\dagger}}),|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
$=$ $(\psi_{j_{1}j_{2}},|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
We continue applying the controlled-$R_{3}$, $R_{4}$ through $R_{n}$ gates and compute our postcondition, each of which adds an an extra bit to the phase of the coefficient of the first |$\ket{1}$. At the end of this procedure, we have the postcondition
$(\psi_{j_{1}j_{2}\cdots j_{n}},|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
Next, we perform a similar procedure on the second qubit. The Hadamard gate puts us in the postcondition
$(\psi_{j_{1}j_{2}\cdots j_{n}},H|j_{2}\rangle\langle j_{2}|H,|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)=(\psi_{j_{1}j_{2}\cdots j_{n}},\psi_{j_{2}},|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
The controlled-$R_{2}$ through $R_{n-1}$ gates yield the predicate
$(\psi_{j_{1}j_{2}\cdots j_{n}},\psi_{j_{2}\cdots j_{n}},|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
We continue in this fashion for each qubit, giving a final predicate
$(\psi_{j_{1}j_{2}\cdots j_{n}},\psi_{j_{2}\cdots j_{n}},\psi_{j_{3}\cdots j_{n}},\cdots,\psi_{j_{n}}).$
It follows the Swap operation between the qubit $i$ and the qubit $n+1-i$ for $1\leq i\leq n$. After applying $SWAP(1,n)$, we obtain
$(\psi_{j_{n}},\psi_{j_{2}\cdots j_{n}},\cdots,\psi_{j_{n-1}j_{n}},\psi_{j_{1}\cdots j_{n}})$
by observing
$\operatorname{supp}({\mathrm{Tr}}_{n}SWAP(1,n)[\psi_{j_{1}\cdots j_{n}}\otimes I_{n}]\cap[I_{1}\otimes\psi_{j_{n}}]SWAP(1,n))$ $=\psi_{j_{n}}$
$\operatorname{supp}({\mathrm{Tr}}_{1}SWAP(1,n)[\psi_{j_{1}\cdots j_{n}}\otimes I_{n}]\cap[I_{1}\otimes\psi_{j_{n}}]SWAP(1,n))$ $=\psi_{j_{1}\cdots j_{n}}.$
After all the swap operations, the postcondition is
$(\psi_{j_{n}},\psi_{j_{n-1}j_{n}},\cdots,\psi_{j_{2}\cdots j_{n}},\psi_{j_{1}\cdots j_{n}})$
The postcondition derived from QAI is an abstract state represented as a tuple of density matrices corresponding to pure quantum states.