### v3

# A Logic for Approximate Quantitative Reasoning about Quantum Circuits

Abstract.
Reasoning about quantum programs remains a fundamental challenge, regardless of the programming model or computational paradigm. Despite extensive research, existing verification techniques are insufficient—even for quantum circuits, a deliberately restricted model that lacks classical control, but still underpins many current quantum algorithms. Many existing formal methods require exponential time and space to represent and manipulate (representations of) assertions and judgments, making them impractical for quantum circuits with many qubits. This paper presents a logic for reasoning in such settings, called SAQR-QC. The logic supports Scalable but Approximate Quantitative Reasoning about Quantum Circuits, whence the name. SAQR-QC has three characteristics: (i) some (deliberate) loss of precision is built into it; (ii) it has a mechanism to help the accumulated loss of precision during a sequence of reasoning steps remain small; and (iii) most importantly, to make reasoning scalable, every reasoning step is local—i.e., it involves just a small number of qubits. We demonstrate the effectiveness of SAQR-QC via two case studies: the verification of GHZ circuits involving non-Clifford gates, and the analysis of quantum phase estimation—a core subroutine in Shor’s factoring algorithm.
copyright: nonejournal: PACMPLjournalvolume: 1journalnumber: CONFarticle: 1journalyear: 2018publicationmonth: 1
1. Introduction
Quantum computing leverages superposition and interference to achieve computational advantages over classical methods for specific tasks. For instance, Shor’s algorithm (Shor94, ) efficiently factors large integers using the Quantum Fourier Transform (QFT) and quantum phase estimation (Kitaev1995, ). The quantum-phase-estimation technique is also a foundational component in quantum-simulation algorithms (Aspuru-Guzik2005, ) and in quantum algorithms for solving linear algebra problems (HHL08, ).
Despite the different computational paradigm, ensuring that quantum programs behave as intended remains a critical challenge. This challenge has spurred significant research into verifying quantum programs using classical computers (BHY19, ; ZYY19, ; zhou2021quantum, ; 10.1145/3453483.3454029, ; Unruh, ; YP21, ; 10.1145/3704873, ; 10.1145/3519939.3523431, ; CACM:CCLLTY25, ). A landmark achievement is Ying’s Quantum Hoare Logic (QHL) (Ying11, ), which extends classical Hoare logic to reason about the partial correctness of quantum programs. QHL uses quantum predicates—semidefinite positive Hermitian operators—as pre- and postconditions, with the quantum Hoare triple (DP06, ) defined as:
(1) $\{A\}P\{B\}\quad\mathrm{iff}\quad\textrm{for all}~\rho,{\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\llbracket P\rrbracket(\rho)).$
Despite this progress, scalable verification remains elusive, even for quantum circuits—a restricted model without classical control that encapsulates many current algorithms. Existing methods struggle with exponential complexity; for instance, while QHL is complete for unitary programs, its application requires computing $U_{P}^{\dagger}BU_{P}$ for a general $n$-qubit unitary $U_{P}$, a task requiring time and space exponential in $n$. This intractability motivates our central question: Can we establish a theoretical foundation for scalable quantitative reasoning about quantum programs that use many qubits? Here, “scalable” means that the size of a proof—the matrices involved and the logical derivation itself—grows polynomially with the number of qubits.
This paper answers this question affirmatively by introducing SAQR-QC (Scalable but Approximate Quantitative Reasoning for Quantum Circuits), a logic for reasoning about programs expressed as quantum circuits. SAQR-QC is designed with three principles: (i) it embraces a deliberate, controlled loss of precision to achieve scalability; (ii) it incorporates mechanisms to keep the accumulated imprecision small across reasoning steps; (iii) all reasoning is local, meaning each step involves only a constant number of qubits, independent of the total system size.
SAQR-QC is inspired by two prior approaches that sit at opposite ends of a spectrum. At one end, QHL—see Equation˜1—offers fully precise quantitative reasoning, but is computationally intractable for large systems. At the other end, Quantum Abstract Interpretation (QAI) (YP21, ) provides a scalable, qualitative framework. A QAI judgment,
$\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}\quad\mathrm{iff}\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathrm{forall}\,\rho,\,\rho\vDash\mathcal{P}~\text{implies}~\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},}}$
uses tuples of local projections $\mathcal{P}$ and $\mathcal{Q}$ as abstract states. QAI enables efficient, local reasoning but lacks the ability to reason about quantitative properties like success probabilities.
SAQR-QC integrates ideas from both to achieve scalable quantitative reasoning. It is important to clarify that SAQR-QC is a logic for manual proof construction, analogous to how Hoare Logic, Separation Logic, and Linear Logic were first developed as conceptual frameworks to facilitate human-driven reasoning. The quest for automation, while a vital future direction, is not the aim of this paper.
A judgment in SAQR-QC has a two-part structure, inspired by the spatial and pure assertions of separation logic (DBLP:conf/fmco/BerdineCO05, ; DBLP:conf/tacas/DistefanoOY06, ):
(2) $\{\mathscr{A}|\mathcal{P}\}C\{\mathscr{B}|\mathcal{Q}\}.$
Here, $\mathcal{P}$ and $\mathcal{Q}$ are QAI-style tuples of local projections, capturing qualitative spatial constraints on the state. $\mathscr{A}$ and $\mathscr{B}$ are tuples of local observables, each acting non-trivially on only a constant number of qubits. These observables track quantitative information, such as success probabilities, in a manner reminiscent of QHL, but restricted to local components. The triple (2) asserts that for any input state $\rho$ satisfying $\mathcal{P}$, the output state $\llbracket C\rrbracket(\rho)$ satisfies $\mathcal{Q}$, and a certain quantitative relationship involving $\mathscr{A}$ and $\mathscr{B}$ holds. More precisely, a SAQR-QC judgment means
(3) $\models^{\rm\textsc{SAQR-QC}}\{\mathscr{A}|\mathcal{P}\}C\{\mathscr{B}|\mathcal{Q}\}\quad\mathrm{iff}\quad\textrm{for all}~\rho,\,\rho\vDash\mathcal{P}~\textrm{implies}~~\begin{aligned} &\textrm{(i)}~\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q},~~\mathrm{and}\\ &\textrm{(ii)}~{\mathrm{Tr}}(M_{\mathscr{A}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),\end{aligned}$
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are derived from $\mathscr{A}$ and $\mathscr{B}$, with the precise definition provided in §˜3.
This two-part structure is not merely stylistic; it is essential under the locality constraint. Without locality, the two components could be combined into a single QHL triple (ZYY19, , Theorems 3.2 and 3.3). However, with the restriction to local projections and observables, this combination is generally impossible, justifying the distinct roles of the QAI-like and QHL-like components. Their interaction—inspired by the reduced product in abstract interpretation (DBLP:conf/popl/CousotC79, , §10.1), as well as the use of abstract interpretation in logic-based tools (Crab, )—helps to control the accumulation of imprecision.
A primary use-case for SAQR-QC is forward reasoning: given a pre-state assertion $\{\mathscr{A}|\mathcal{P}\}$ and a circuit $C$, derive a post-state assertion $\{\mathscr{B}|\mathcal{Q}\}$. The derived post-condition might not be the strongest possible, but aims to be strong enough to extract useful quantitative information, such as the success probability of a computation in the BQP model. [^1] BQP is the class of decision problems solvable by a quantum computer in polynomial time with bounded error, analogous to how P is defined for classical deterministic computers.
As with any manual logic, constructing a SAQR-QC proof requires insight, particularly in choosing the appropriate local projections and observables. The logic provides the framework, but the user must supply the clever insights—the “eureka” steps—to define the predicates that make the verification possible. The case studies in §˜4 and §˜5 exemplify this principle, demonstrating how specific, carefully chosen predicates yield sharp results for non-trivial quantum circuits.
Contributions
This paper makes the following contributions:
  • We present SAQR-QC, a logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits. SAQR-QC integrates quantitative, QHL-like reasoning with the scalable, local reasoning of QAI.
  • We provide a formal foundation for SAQR-QC, defining its assertions and proof rules. The design ensures that all reasoning steps are local, leading to proofs whose size is polynomial in the number of qubits and gates.
  • We demonstrate the logic’s utility and expressiveness through two detailed case studies:
    • Verification of GHZ circuits involving non-Clifford gates: We derive a judgment that characterizes the output state precisely in every aspect except a relative phase factor.
    • Analysis of Quantum Phase Estimation (QPE): We derive a judgment showing that for any constant $k$, the QPE algorithm provides the best estimate of the last $k$ bits of the phase, with probability at least $4/\pi^{2}$. A key step is a novel, lossless local-reasoning method for the Quantum Fourier Transform using QAI. To the best of our knowledge, no prior approach has achieved such a result with a proof that scales polynomially with the system size.
  • We clarify the relationship between SAQR-QC and prior work, particularly Zhou et al. (ZYY19, ), and discuss the principles for constructing effective predicates for use in our framework.
In summary, this work takes a foundational step towards scalable reasoning about quantum computations by defining a logic that is both quantitative and tractable, and by demonstrating its power on representative quantum algorithms.
Organization of the paper
§˜2 presents background material about quantum computing, as well as the basics of QAI. §˜3 defines local observables as predicates, and formalizes the judgments of SAQR-QC, which integrate local observables and QAI. We employ SAQR-QC to reason about the general GHZ circuit (§˜4) and quantum phase estimation (§˜5). §˜6 discusses related work. §˜7 concludes. Some proofs and derivations are available in the paper’s appendices, which are submitted as Supplementary Material.
2. Background & Notation
To make the paper self-contained, this section reviews basic concepts, notation, and properties of quantum computing, following standard references such as (NI11, ; YP21, ; Ying11, ), as well as the basics of QAI. Readers familiar with these topics may wish to skip to §˜3.
2.1. Preliminaries
[…]
Linear operators on $d$-dimensional complex vector spaces are represented by $d\times d$ matrices $\mathbb{C}^{d\times d}$. The identity is $I$, and the conjugate transpose of $A$ is $A^{\dagger}=(A^{T})^{*}$. An operator is Hermitian if $A=A^{\dagger}$ and positive semi-definite if all eigenvalues are non-negative. The trace, ${\mathrm{Tr}}(A)$, is the sum of the diagonal entries: ${\mathrm{Tr}}(A)=\sum_{i}A_{ii}$.
The Löwner order on Hermitian matrices, $A\leq B$ if BAB-A is positive semidefinite, is fundamental for comparing states and operators in quantum mechanics.
2.2. Quantum States
[…]
2.3. Reduced Density Matrices
[…]
For an $n$-partite system and $s\subseteq[n]$, the reduced density matrix is
$\rho_{s}={\mathrm{Tr}}_{[n]\setminus s}(\rho),$
with ${\mathrm{Tr}}_{[n]\setminus s}$ tracing out all qubits not in $s$. The partial trace preserves positive semi-definiteness (NI11, ).
2.4. Unitary Operations
[…]
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\qquad X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\qquad Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}\qquad Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\qquad\textit{SWAP}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix}$
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\qquad T=\begin{pmatrix}1&0\\ 0&e^{i\pi/4}\end{pmatrix}\qquad R_{m}={\begin{pmatrix}1&0\\ 0&e^{2\pi i/2^{m}}\end{pmatrix}}\qquad\textit{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$
2.5. Observables
Quantum observables represent measurable quantities, such as position, energy, or spin, and are modeled by Hermitian operators $O$ with $O^{\dagger}=O$, ensuring real outcomes. In contexts like Quantum Hoare Logic (QHL), observables are often restricted to $0\leq O\leq I$ in the Löwner order (§˜2.1), meaning both $O$ and IOI-O are positive semidefinite. These bounded observables serve as predicates or quantum effects, capturing partial truth values in program verification.
2.6. Quantum Circuits and Semantics
[…]
Each gate is lifted to an $n$-qubit unitary by tensoring with identities. For a single-qubit gate $U$ acting on qubit $i$, $U\otimes I_{[n]\setminus\{i\}}:=(\otimes_{k>i}I)\otimes U\otimes(\otimes_{0<j<i}I)$, and similarly for two-qubit gates $U_{F_{\ell}}$ on $F_{\ell}=\{i,j\}$, $U\otimes I_{[n]\setminus\{i,j\}}$, with placement determined by the qubit indices. For clarity, we describe semantics assuming 2-qubit gates, though all results extend to gates acting on up to $m$ qubits for any constant $m$.
Definition 2.1 (Syntax).
The syntax of quantum programs is given by
$C::=\mathbf{skip}\mid\bar{q}:=U[\bar{q}]\mid C_{1};C_{2}$
We write $\llbracket C\rrbracket$ to denote the semantics of a quantum program $C$. If $C$ represents a unitary transformation $U_{C}$, then for any input density matrix $\rho$, its semantics is given by
$\llbracket C\rrbracket(\rho):=U_{C}\rho U_{C}^{\dagger}.$
We also define the dual action on observable $A$ as:
$\llbracket C\rrbracket^{*}(A):=U_{C}^{\dagger}AU_{C}.$
We define $\mathcal{U}:=\lambda x.UxU^{\dagger}$ to denote the quantum operation on reduced density matrices that corresponds to the unitary matrix $U$ (which operates on quantum states).
2.7. Projections
An orthogonal projection satisfies $P=P^{\dagger}=P^{2}$, a stronger condition than the classical $P=P^{2}$. We simply call such matrices projections. Each projection $P$ corresponds to a subspace $S_{P}=\{v\mid Pv=v\}$, and we use “projection” and “subspace” interchangeably. This induces a partial order: $P\subseteq Q$ iff $S_{P}\subseteq S_{Q}$. For example, $|00\rangle\langle 00|+|11\rangle\langle 11|$ is a rank-2 projection onto a 2D subspace of $\mathbb{C}^{4}$.
Projections are positive semidefinite. The support of a positive semidefinite matrix $A$, $\operatorname{supp}(A)$, is the span of eigenvectors with nonzero eigenvalues. A density matrix $\rho$ satisfies $P$, written $\rho\vDash P$, if $\operatorname{supp}(\rho)\subseteq P$, equivalently $P\rho=\rho$ (BvN36, ).
2.8. Lemmas
Our development of SAQR-QC relies on three fundamental operations on operators:
  • Löwner order of operators (denoted $A\leq B$),
  • Partial trace and trace operators (denoted ${\mathrm{Tr}}_{s}$ and $\mathrm{Tr}$}, where ${\mathrm{Tr}}_{s}$ traces out subsystem $s$),
  • Expansion of an operator via tensor product (denoted $A_{s}\otimes I_{[n]\setminus s}$, where the operator $A_{s}$ acts on subsystem $s$ and is expanded to the full system).
The following lemmas summarize key algebraic relationships among these operations, forming the basis for the correctness of our SAQR-QC framework. Unless stated otherwise, let $s\subseteq[n]$; $P$ be a projection on an $n$-qubit system; $A,B$ be positive semidefinite matrices; $E$ a matrix; and $\rho$ a quantum state. The proof of ˜2.1 is given in Appendix˜A. The other lemmas can be proven using the definition of trace, partial trace, and support.
The relevance of these lemmas to SAQR-QC is that (i) a positive semidefinite matrix $A$ serves as a predicate on a density matrix $\rho$ via the expectation value ${\mathrm{Tr}}(A\rho)$; and (ii) a subclass of such predicates with good locality properties are ones defined by a positive semidefinite matrix $A_{s}$ that only acts on qubit set $s\subseteq[n]$, producing the expectation value ${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right)$.
Lemma 2.1.
Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$, ${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right)$.
Lemma 2.2.
For two square matrices $B$ and $E$, ${\mathrm{Tr}}(BE)={\mathrm{Tr}}(EB)$.
Lemma 2.3.
For $A,B\geq 0$, ${\mathrm{Tr}}(AB)\geq 0$.
Lemma 2.4.
If $A\leq B$, then ${\mathrm{Tr}}(A\rho)\leq{\mathrm{Tr}}(B\rho)$ for any density operator $\rho$, and in particular ${\mathrm{Tr}}(A)\leq{\mathrm{Tr}}(B)$.
Lemma 2.5.
For a density matrix $\rho$ and a projection $P$, we have
${\mathrm{Tr}}(P\rho)=1\quad\Leftrightarrow\quad\operatorname{supp}(\rho)\subseteq P\quad\Leftrightarrow\quad\rho\vDash P.$
2.9. Qualitative Predicates for Local Reasoning: Quantum Abstract Interpretation
The qualitative predicates used in QAI express logical properties of quantum states, e.g., whether a subsystem lies in a subspace. Represented as tuples of local projectors, they support scalable local reasoning, inferring global behavior from partial views. QAI (YP21, ) propagates these predicates through circuits using small-subsystem projectors, avoiding exponential full-state analysis.
This section reviews local projective predicates, formalizes their semantics, and describes their transformation under unitaries. Soundness is established via support-based semantics and partial trace, forming the basis for abstract reasoning about quantum programs.
Definition 2.1 ((YP21, )).
A tuple $(P_{s_{1}},\cdots,P_{s_{m}})$ is called a projective predicate if each $P_{s_{i}}$ is a projection, i.e., $P_{s_{i}}^{2}=P_{s_{i}}$. We use $\mathcal{P}$ (or $\mathcal{Q},\mathcal{R}$) to denote projective predicates. In particular, we write $\mathcal{I}:=(I_{s_{1}},\cdots,I_{s_{m}})$ to represent the identity predicate.
Remark. Projective predicates are also referred to as abstract states.
Definition 2.2 ((YP21, )).
A state $\rho$ satisfies a projective predicate $\mathcal{P}=(P_{s_{1}},\dots,P_{s_{m}})$, denoted by
$\rho\vDash^{QAI}\mathcal{P},$
if for all $1\leq i\leq m$, $P_{s_{i}}\rho_{s_{i}}=\rho_{s_{i}}$, i.e., $\rho_{s_{i}}\vDash P_{s_{i}}$. Equivalently, $\rho\vDash\gamma(\mathcal{P})$, where
$\gamma(\mathcal{P}):=\bigcap_{i}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Given a quantum circuit $\mathbf{C}$ and a state $\rho\vDash\mathcal{P}$, the abstract-interpretation method presented in (YP21, ) constructs a predicate $\mathcal{Q}$ such that the post-state $\llbracket\mathbf{C}\rrbracket(\rho)$ satisfies $\mathcal{Q}$, denoted as
$\vDash^{QAI}\{\mathcal{P}\}{C}\{\mathcal{Q}\}$
Theorem 2.1 ((YP21, )).
Let $U_{F}$ be a unitary gate applied to the qubit set $s(F)$, and let $\mathcal{P}=(P_{s_{1}},\cdots,P_{s_{m}})$ be a projective predicate. For each $s_{i}$, define
[…]
Then
$\rho\vDash^{QAI}\mathcal{P}\quad\Rightarrow\quad U_{F}\rho U_{F}^{\dagger}\vDash^{QAI}U^{\sharp}(\mathcal{P}).$
Theorem 2.2 ((YP21, )).
[…]
3. Correctness Formulas and the Logical System SAQR-QC
In this section, we introduce a class of predicates that can be used for quantitative local reasoning (§˜3.1). We then sketch a strawman approach that incorporates these predicates in a logic similar to Quantum Hoare Logic (QHL) (§˜3.2). Via a simple example, we demonstrate that this strawman approach can lead to a significant loss of precision in reasoning. To address this limitation, we integrate the QAI technique into the framework, leading to the formal definition of the judgments used in SAQR-QC (§˜3.3), a theorem about how SAQR-QC judgments relate to judgments in QAI (§˜3.4), a systematic presentation of the inference rules of SAQR-QC (§˜3.5), and an approximation strategy that addresses how to work with the one inference rule of SAQR-QC that poses a challenge to scalability (§˜3.6).
3.1. Quantitative Local Reasoning via Generalized Predicates
Reasoning about full quantum states quickly becomes infeasible due to the exponential growth of the state space with the number of qubits. To address this issue, we want an approach akin to QAI (YP21, ), which uses tuples of projections as predicates that capture properties of small subsystems. However, QAI itself is unsatisfactory because QAI can only perform qualitative local reasoning.
To enable quantitative local reasoning, we focus on reduced density matrices, which capture the behavior of local subsystems. Inspired by Quantum Hoare Logic (QHL)—where a global positive semidefinite matrix $A$ serves as a predicate describing properties of a state $\rho$ via the expectation value ${\mathrm{Tr}}(A\rho)$—we introduce the notion of local observables to reason about subsystem states $\rho_{s}$.
Definition 3.1.
A local observable over a qubit set $s\subseteq[n]$ is a positive semidefinite operator $A_{s}$ satisfying $0\leq A_{s}\leq I_{s}$, acting nontrivially only on the qubits in $s$.
According to Lemma 2.1, ${\mathrm{Tr}}(A_{s}\rho_{s})={\mathrm{Tr}}(A\rho)$ with $A:=A_{s}\otimes I_{[n]\setminus s}$. Local observables thus form a natural subclass of predicates in QHL, and we use them to monitor reduced density matrices via the inner product ${\mathrm{Tr}}(\cdot)$, in direct analogy with QHL; here, “monitoring $\rho_{s}$” means computing the expectation value of local observable $A_{s}$ with respect to $\rho_{s}$—i.e., computing ${\mathrm{Tr}}(A_{s}\rho_{s})$.
To capture richer information, we track not a single $\rho_{s}$ but a tuple of reduced density matrices $(\rho_{s_{1}},\dots,\rho_{s_{m}})$, where each $s_{i}\subseteq[n]$ is a small subset of bounded size; we also refer to these as low-dimensional marginals. This representation (i) provides significantly deeper insight into the system’s local structure, and (ii) remains tractable, because it requires only linear resources.
The idea is to define a quantum predicate for each $s_{i}$, represented by an observable $A_{s_{i}}$, focusing solely on the state of the quantum registers in $s_{i}$.
Definition 3.2.
A predicate over an $n$-qubit space and $S=(s_{1},\dots,s_{m})$ is a tuple of local observables
$(A_{s_{1}},\dots,A_{s_{m}})\quad\text{with}\quad 0\leq A_{s_{i}}\leq I_{s_{i}},$
each acting nontrivially only on $s_{i}$. Predicates are denoted by $\mathscr{A}$ (or $\mathscr{B},\mathscr{D}$); $m$ is the predicate’s size.
Each local observable acts nontrivally on just a few qubits, and captures the expectation value of a measurable quantity, such as the success probability, fidelity, or entanglement, providing a direct link between semantics and operational outcomes. By tracing their evolution under unitaries and partial traces, we obtain a scalable, compositional framework for quantum program verification, supporting precise analysis without reconstructing the full state.
These generalized predicates track quantitative aspects of quantum programs—such as probabilities and expectation values—by monitoring reduced density matrices over small, constant-size subsystems. This abstraction underpins scalable, compositional reasoning about quantum computations via low-dimensional, information-preserving summaries.
Definition 3.3.
For a predicate $\mathscr{A}=(A_{s_{1}},\cdots,A_{s_{m}})$, its matrix representation $M_{\mathscr{A}}$ is defined as:
$\small M_{\mathscr{A}}=\sum_{i=1}^{m}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
This definition tracks changes in a tuple of reduced density matrices via linear functionals, enabling scalable, compositional quantitative reasoning. By Lemma 2.1, the trace ${\mathrm{Tr}}(M_{\mathscr{A}}\rho)$ depends only on the reduced states $\rho_{s_{i}}$:
(4) $\small{\mathrm{Tr}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}\Big(\sum_{i=1}^{m}(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho\Big)=\sum_{i=1}^{m}{\mathrm{Tr}}[(A_{s_{i}}\otimes I_{[n]\setminus s_{i}})\rho]=\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}\rho_{s_{i}}).$
Definition 3.4.
The domain of a predicate $(A_{s_{1}},\dots,A_{s_{m}})$ is $\mathrm{dom}(A_{s_{1}},\dots,A_{s_{m}})=(s_{1},\dots,s_{m})$.
We are also interested in the order relation between predicates.
Definition 3.5.
Given predicates $(A_{s_{1}},\cdots,A_{s_{m}})$ and $(A_{s_{1}}^{\prime},\cdots,A_{s_{m}}^{\prime})$ over the same domains, define
$(A_{s_{1}},\cdots,A_{s_{m}})\sqsubseteq(A_{s_{1}}^{\prime},\cdots,A_{s_{m}}^{\prime})\quad\text{if and only if}\quad A_{s_{i}}\leq A_{s_{i}}^{\prime}\text{ for all }i.$
Lemma 3.1.
The matrix representation of predicates is monotonic with respect to this ordering—i.e.,
$\mathscr{A}\sqsubseteq\mathscr{B}\quad\Rightarrow\quad M_{\mathscr{A}}\leq M_{\mathscr{B}}.$
3.2. A First Attempt at Defining Judgments
This section presents an initial attempt for quantitative local reasoning in quantum programs using tuples of local observables.
In QHL (Ying11, ), the unitary rule gives the weakest precondition for a postcondition:
$\{U_{P}^{\dagger}BU_{P}\}\ P\ \{B\}.$
Let $S=(s_{1},\dots,s_{m})$ with $s_{i}\subseteq[n]$, and $\mathscr{B}=(B_{s_{1}},\dots,B_{s_{m}})$ with $0\leq B_{s_{i}}\leq I_{s_{i}}$; then
$M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Definition 3.6.
Given a fixed domain $S=(s_{1},\dots,s_{m})$ and a quantum program $\mathbf{C}=\lambda\rho.\,U\rho U^{\dagger}$ on a density matrix, we call $\mathscr{A}=(A_{s_{1}},\dots,A_{s_{m}})$ a local precondition of $\mathscr{B}=(B_{s_{1}},\dots,B_{s_{m}})$ if
$\{M_{\mathscr{A}}\}\,C\,\{M_{\mathscr{B}}\},$
where such a judgment means
$\models\{M_{\mathscr{A}}\}\,C\,\{M_{\mathscr{B}}\}\quad\mathrm{iff}\quad\textrm{for all}~\rho,\,~{\mathrm{Tr}}(M_{\mathscr{A}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),$
with $M_{\mathscr{A}}=\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ and $M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}$.
We could use a QHL-like proof rule for a unitary, namely
[…]
To see that $M_{\mathscr{A}}\leq U^{\dagger}M_{\mathscr{B}}U$ leads to a valid QHL judgement—without requiring that the observables are bounded above by the identity—we observe that
(6) ${\mathrm{Tr}}\left(M_{\mathscr{A}}\rho\right)\leq{\mathrm{Tr}}\left(U^{\dagger}M_{\mathscr{B}}U\rho\right)={\mathrm{Tr}}\left(M_{\mathscr{B}}U\rho U^{\dagger}\right)={\mathrm{Tr}}\left(M_{\mathscr{B}}\,\llbracket\mathbf{C}\rrbracket(\rho)\right),$
where the inequality follows from ˜2.4 applied to the premise of Rule 5, and the first equality uses ˜2.2. ˜3.6 tells us that the conclusion of Rule 5 holds.
This definition naturally gives rise to a correctness judgment of the form $\{\mathscr{A}\}\ \mathbf{C}\ \{\mathscr{B}\}$ for quantum circuits. In addition to enabling backward reasoning, Rule 5 with ˜3.6 also supports forward reasoning by changing the premise of Rule 5 to
$UM_{\mathscr{A}}U^{\dagger}\leq M_{\mathscr{B}}.$
Unfortunately, it is not possible, in general, to use either version of Rule 5 algorithmically to compute the weakest precondition and the strongest postcondition, respectively. The issue is one of expressibility: can the answer be decomposed according to the chosen scheme $S=(s_{1},\ldots,s_{m})$? In general, this structural constraint of our candidate logic presents an obstacle. For instance, even when $U_{F}$ is a two-qubit unitary, the transformed observable
$U_{F}^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U_{F}$
may not admit a decomposition of the form $\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ for any choice of local observables $A_{s_{i}}$. That is, the following equality may have no solution for local predicates $\mathscr{A}$, for some given $\mathscr{B}$.
(7) $M_{\mathscr{A}}=U_{F}^{\dagger}M_{\mathscr{B}}U_{F}$
Example 3.1.
To understand this limitation concretely, consider a circuit in which a CNOT gate is applied to the input state $|0\ket$$\ket{0}$, as shown in Figure 1, and let us reason about the circuit $\mathbf{C}$ using Equation˜7. We choose $(s_{1},s_{2})=(\{q_{1}\},\{q_{2}\})$ and define the postcondition to be $M_{\mathscr{B}}=|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$. Observation: The following calculation will demonstrate that, in local reasoning, it is not always possible to express the weakest precondition.
We will show that there is no $A_{1},A_{2}$ such that
$U_{\mathbf{C}}^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U_{\mathbf{C}}$ $=|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|$
$=A_{1}\otimes I+I\otimes A_{2}$
Figure 1. CNOT circuit $\mathbf{C}$.
We see that the left-hand side is orthogonal to $|1\rangle\langle 1|\otimes|0\rangle\langle 0|$. Therefore, we know that
${\mathrm{Tr}}[(|1\rangle\langle 1|\otimes|0\rangle\langle 0|)(A_{1}\otimes I+I\otimes A_{2})]=0$
$\Longrightarrow$ ${\mathrm{Tr}}[(|1\rangle\langle 1|A_{1})\otimes|0\rangle\langle 0|]+{\mathrm{Tr}}[|1\rangle\langle 1|\otimes(|0\rangle\langle 0|A_{2})]=0$
[…]
$\Longrightarrow$ $A_{1}=\lambda_{1}|0\rangle\langle 0|\quad\textrm{{and}}\quad A_{2}=\lambda_{2}|1\rangle\langle 1|$
$\Longrightarrow$ $A_{1}\otimes I+I\otimes A_{2}=\lambda_{1}|0\rangle\langle 0|\otimes I+\lambda_{2}I\otimes|1\rangle\langle 1|\neq|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|.$
[…]
(8) $\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}.$
To establish that Rule 5 cannot prove that Assertion 8 holds, we compute
$\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}$
$\Longleftrightarrow$ $\mathit{CNOT}(|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|)\mathit{CNOT}^{{\dagger}}\leq|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$
$\Longleftrightarrow$ $|0\rangle\langle 0|\otimes I+|00\rangle\langle 00|+|11\rangle\langle 11|\leq|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$
$\Longleftrightarrow$ $|00\rangle\langle 00|+|11\rangle\langle 11|\leq I\otimes|0\rangle\langle 0|$
Because the final inequality fails, in our strawman logic based on tuples of local observables and Rule 5, Assertion 8 is unprovable; i.e.,
$\not\vdash\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}.$
3.3. Quantitative Judgments and Validity
[…]
Definition 3.7.
A judgment is a triple of the form
$\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}$
for program $\mathbf{C}$, general predicates $\mathscr{A}$ and $\mathscr{B}$, and projective predicates $\mathcal{P}$ and $\mathcal{Q}$.
Definition 3.8 (Validity).
The judgment $\{\mathscr{A}|\mathcal{P}\}\,\mathbf{C}\,\{\mathscr{B}|\mathcal{Q}\}$ in ˜3.7 holds if $\forall\rho\vDash^{QAI}\mathcal{P}$,
(9) $\llbracket\mathbf{C}\rrbracket(\rho)$ $\vDash^{QAI}\mathcal{Q},$
(10) ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ $\leq{{\mathrm{Tr}}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),$
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are the matrix representations of $\mathscr{A}$ and $\mathscr{B}$, as defined in ˜$3.$
Consider the special case $\mathcal{P}=\mathcal{Q}=\mathcal{I}$ with $\mathcal{I}:=(I_{s_{1}},\cdots,I_{s_{m}})$, we will write
$\{\mathscr{A}|\mathcal{I}\}\mathbf{C}\{\mathscr{B}|\mathcal{I}\}=\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}$
in which case this definition simplifies to ˜3.6.
We have that ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ only depends on the reduced density matrices of $\rho$ with respect to the systems $\set{s_{i}}$, as we know from Equation˜4,
${\mathrm{Tr}}(M_{\mathscr{A}}\rho)=\sum_{i=1}^{m}{\mathrm{Tr}}\left(A_{s_{i}}\rho_{s_{i}}\right).$
In other words, ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ only depends on the reduced density matrices of $\rho$ on the systems $\set{s_{i}}$. Intuitively, this judgment tracks a linear function of the tuple of reduced density matrices to enable quantitative reasoning.
Moreover, $\rho_{s_{i}}\vDash P_{s_{i}}$ implies that $\rho_{s_{i}}=P_{s_{i}}\rho_{s_{i}}P_{s_{i}}$. This property implies
$\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}\rho_{s_{i}})=\sum_{i=1}^{m}{\mathrm{Tr}}(A_{s_{i}}P_{s_{i}}\rho_{s_{i}}P_{s_{i}})=\sum_{i=1}^{m}{\mathrm{Tr}}(P_{s_{i}}A_{s_{i}}P_{s_{i}}\rho_{s_{i}}).$
where the last step is due to ˜$2.$
We observe that ${\mathrm{Tr}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}(M_{\mathscr{A}^{\prime}}\rho)$ with $\mathscr{A}^{\prime}=(P_{s_{1}}A_{s_{1}}P_{s_{1}},\dots,P_{s_{m}}A_{s_{m}}P_{s_{m}})$. Then
$\{\mathscr{A}^{\prime}|\mathcal{P}\}\,\mathbf{skip}\,\{\mathscr{A}|\mathcal{P}\}\quad\text{and}\quad\{\mathscr{A}|\mathcal{P}\}\,\mathbf{skip}\,\{\mathscr{A}^{\prime}|\mathcal{P}\}.$
Thus, the qualitative insights from quantum abstract interpretation refine the reasoning process by yielding “stronger" postconditions and “weaker" preconditions. [^2] Here, “stronger” and “weaker” simply refer to the relation $\mathrm{tr}(A_{s_{i}})\geq\mathrm{tr}\big((P_{s_{i}}P_{s_{i}})A_{s_{i}}\big)\geq\mathrm{tr}\big(P_{s_{i}}A_{s_{i}}P_{s_{i}}\big),$ as established by Lemma 2.2. They do not indicate any ordering relationship between $\mathscr{A}$ and $\mathscr{A}^{\prime}$.
3.4. Reduction
We now consider the relationship between the correctness $\models^{\rm QAI}$ in QAI and $\models$ in ˜3.8.
Theorem 3.1 (Reduction Principle).
Let the quantum program be a quantum circuit, and consider its behavior with respect to any input state. For any projective predicates $\mathcal{P}=(P_{s_{1}},\cdots,P_{s_{m}})$ and $\mathcal{Q}=(Q_{s_{1}},\cdots,Q_{s_{m}})$, $\mathcal{P}$ and $\mathcal{Q}$ can be regarded as the observables $\mathscr{P}:=(P_{s_{1}},\cdots,P_{s_{m}})$ and $\mathscr{Q}=(Q_{s_{1}},\cdots,Q_{s_{m}})$, respectively. Then we have the property
If $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ in the sense of ˜3.8, then $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$.
We defer the proof to the appendix.
3.5. Logical System with Soundness
The inference rules for program constructs in SAQR-QC are given in Figure˜2.
$\begin{split}&\textsc{Skip}\quad\ \frac{}{\{\mathscr{A}|\mathcal{P}\}\mathbf{Skip}\{\mathscr{A}|\mathcal{P}\}}\\[2.84544pt] &\textsc{Unit}\quad\ \frac{\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})}{\{\mathscr{A}|\mathcal{P}\}\bar{q}:=U_{F}\left[\bar{q}\right]\{\mathscr{B}|U_{F}^{\sharp}(\mathcal{P})\}}\ \ \ \ \ \ \\[2.84544pt] &\textsc{Seq}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1}\{\mathscr{D}|\mathcal{R}\}\ \ \ \ \ \ \{\mathscr{D}|\mathcal{R}\}\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1};\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}\ \ \ \ \ \ \\[2.84544pt] &\textsc{Con}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\},\ \ \mathscr{D}\sqsubseteq\mathscr{A},\ \mathscr{B}\sqsubseteq\mathscr{E},\ \mathcal{R}\sqsubseteq\mathcal{P},\ \mathcal{Q}\sqsubseteq\mathcal{T}}{\{\mathscr{D}|\mathcal{R}\}\mathbf{C}\{\mathscr{E}|\mathcal{T}\}}\end{split}$
Figure 2. Inference rules for program constructs in SAQR-QC. We can use the proof rules for both forward reasoning or backward reasoning.
Theorem 3.2 (Soundness).
The proof system in Figure˜2 is sound. That is, for quantum program $\mathbf{C}$, $\vdash\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}\ {\rm implies}\ \models\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}.$
We defer the proof to the appendix.
3.6. Scalable Application of the Unit Rule
Among the proof rules in Figure 2, only the rule Unit poses a significant challenge to scalability. Specifically, the condition
(11) $\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\gamma(\mathcal{P})$
requires computing the projector $\gamma(\mathcal{P})$, which becomes intractable for systems with many qubits. That is, given a postcondition $\mathscr{B}$ and a set of projective predicates $\mathcal{P}$, there is no scalable method for synthesizing a precondition $\mathscr{A}$ satisfying this inequality.
To address this issue, we introduce a compositional approximation strategy. The key idea is to construct inequalities over high-dimensional systems by composing inequalities over smaller subsystems. These approximations trade some precision for tractability, but preserve overall soundness and enable reasoning about large quantum programs.
Warm-up: Ignoring $\gamma(\mathcal{P})$
As a first step—which, although imprecise (see §˜3.2), may still yield useful results (as in §˜4)—we omit the projection $\gamma(\mathcal{P})$. This approximation gives rise to a simple linear inequality:
(12) $M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\quad\Longleftrightarrow\quad\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\leq U^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U.$
To improve scalability, we partition the index set $\{1,\dots,m\}$ into disjoint subsets $T_{1},\dots,T_{k}$, where each $T_{j}$ corresponds to a group of local predicates. We establish soundness via localized reasoning:
(13) $\forall j.\quad\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\leq U^{\dagger}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\quad\Longrightarrow\quad M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}}).$
The forward form of this inequality is:
(14) $\forall j.\quad U\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U^{\dagger}\leq\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Because each inequality involves a small subsystem, constraints can be efficiently discharged using symbolic solvers or semidefinite programming, depending on the structure of $A_{s_{i}}$ and $B_{s_{i}}$.
A More Precise Approach: Guarded Inequalities
The warm-up approach may lose critical precision due to ignoring $\gamma(\mathcal{P})$. To refine this, we incorporate the projection information in a compositional way. Let $\mathcal{P}=\{P_{s_{1}},\dots,P_{s_{m}}\}$, and again partition the index set into disjoint subsets $T_{1},\dots,T_{k}$. For each $T_{j}$, we define a local over-approximation of the global projector:
(15) $\gamma(\mathcal{P})\subseteq P_{j}:=\bigcap_{i\in T_{j}}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
We can write guarded versions of the local inequalities:
(16) $\forall j.\quad P_{j}\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)P_{j}\leq P_{j}U^{\dagger}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)UP_{j}.$
These inequalities are more precise than those in (13), as they account for the structure imposed by $\mathcal{P}$, while still avoiding the need to compute the full projector $\gamma(\mathcal{P})$.
Summary of Reasoning Tools
We thus obtain two practical tools for scalable reasoning under the Unit rule:
  • Unguarded inequalities: (13) and (14), which enable fast reasoning but may be imprecise;
  • Guarded inequalities: (16), which retain more logical structure from $\gamma(\mathcal{P})$ while remaining tractable.
Finally, in Appendix˜D, we prove that the guarded inequalities imply the full condition (11), thereby validating the correctness of our compositional approximation framework.
Remark: The methodology presented in this section is general. For a given circuit, useful predicates can typically be constructed straightforwardly because each gate in a quantum circuit is represented by a local unitary operator that typically acts on at most two qubits. In each reasoning step, we analyze a single constant-qubit unitary—typically a one- or two-qubit gate—and extract the predicate $B_{s_{i}}$ over its qubits, leaving the remainder untouched. This approach is sound because the unitary does not affect the reduced density matrices of qubits outside its support. The convenience and effectiveness of this approach are illustrated in the examples throughout the paper.
4. Quantitative Reasoning of general GHZ circuit
We analyze a generalized GHZ circuit where half the gates are arbitrary single-qubit unitaries, producing a highly entangled, densely parameterized state beyond the scope of classical simulation methods such as the Gottesman–Knill theorem (Aaronson_2004, ). Using the scalable techniques developed herein, our analysis proceeds by first applying the “warm-up” method from §˜3.6 (Equation˜13), which enables compositional and projection-free reasoning over subsystems, thereby supporting tractable approximation of the circuit’s complex output state.
Figure 3. GHZ circuit with unitaries $U_{i}$.
In the second phase, we leverage Quantum Abstract Interpretation (QAI), employing the Assert rule from ˜2.2 to confine the output state within a two-dimensional subspace. This critical dimensionality reduction facilitates precise amplitude approximations within our quantitative reasoning framework. Strikingly, the resulting characterization matches the exact wavefunction up to a phase, underscoring that reasoning in SAQR-QC is not merely procedural but requires judicious predicate selection aligned with the circuit’s structure—demonstrating both precision and flexibility when applied with insight.
We select the domain $(\{1,2\},\{2,3\},\cdots,\{n-1,n\})$ for reasoning about Figure 3.
STEP 1. We first select the precondition to be $\{\mathscr{A}|\mathcal{P}\}$
$\mathscr{A}$ $=(A_{1,2},A_{2,3},\cdots,A_{n-1,n})=(|++\rangle\langle++|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|)$
$\mathcal{P}$ $=(P_{1,2},P_{2,3},\cdots,P_{n-1,n})=(|00\rangle\langle 00|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|).$
One can verify that the initial state $|0\cdots 0\rangle\langle 0\cdots 0|\vDash\mathcal{P}$. We now use our proof rules to compute a postcondition for the GHZ circuit, given the precondition $\{\mathscr{A}|\mathcal{P}\}$.
After the first $H$ gate, we can use inequality 12 to derive the following:
$\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus\{i,i+1\}}\leq H^{{\dagger}}\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}H$
$\Longleftrightarrow$ $\sum_{i}H{A_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}H^{{\dagger}}\leq\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}$
$\Longleftrightarrow$ $H{A_{1,2}}H^{{\dagger}}\otimes I_{[n]\setminus\{1,2\}}+\sum_{i>1}{A_{i,i+1}}\otimes HI_{[n]\setminus\{i,i+1\}}H^{{\dagger}}\leq\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}$
$\Longleftrightarrow$ $H{A_{1,2}}H^{{\dagger}}\otimes I_{[n]\setminus\{1,2\}}+\sum_{i>1}{A_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}\leq\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}.$
The final inequality above matches inequality 14 on a term-by-term basis. We can satisfy inequality 14 when each of the following single-term inequalities hold:
$H{A_{1,2}}H^{{\dagger}}\otimes I_{[n]\setminus\{1,2\}}\leq B_{1,2}\otimes I_{[n]\setminus\{1,2\}}\qquad\quad A_{i,i+1}\otimes I_{[n]\setminus\{i,i+1\}}\leq B_{i,i+1}\otimes I_{[n]\setminus\{i,i+1\}}\ \mathrm{for}\ i>1.$
Each inequality can be realized as an equality, yielding the postcondition $\mathscr{B}\mid\mathcal{Q}$},
$\mathscr{B}=(B_{1,2},B_{2,3},\cdots,B_{n-1,n})=(HA_{1,2}H^{{\dagger}},A_{2,3},\cdots,A_{n-1,n})=(|0+\rangle\langle 0+|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|)$
$\mathcal{Q}=(Q_{1,2},Q_{2,3},\cdots,Q_{n-1,n})=(HP_{1,2}H^{{\dagger}},P_{2,3},\cdots,P_{n-1,n})=(|+0\rangle\langle+0|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|)$
After the first CNOT gate has been applied to qubits $q_{1}q_{2}$, we use the Unit Rule to obtain a postcondition $\{\mathscr{C}|\mathcal{R}\}:=\{(C_{1,2},C_{2,3},C_{3,4}\cdots,C_{n-1,n})|(R_{1,2},R_{2,3},R_{3,4}\cdots,R_{n-1,n})\}$. We use QAI to compute $R_{1,2}=|00\rangle\langle 00|+|11\rangle\langle 11|$, $R_{2,3}=|00\rangle\langle 00|+|10\rangle\langle 10|$, and $R_{i,i+1}=|00\rangle\langle 00|$ for $i>2$.
What is left to determine are suitable values for $C_{1,2}$ and $C_{2,3}$. According to the Unit Rule, $\mathscr{C}=(C_{1,2},C_{2,3},C_{3,4}\cdots,C_{n-1,n})$ satisfies
$\text{CNOT}_{1,2}(\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
$\Longleftrightarrow$ $\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}+\sum_{i>2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
$\Longleftrightarrow$ $\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}+\sum_{i>2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
At this point, we can satisfy inequality 14 by satisfying the following smaller inequalities:
$\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i\leq 2,\ \ }I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1},$
$\mathrm{and}\ \ (I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\leq(I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1})\ \mathrm{for}\ i>2.$
We can choose $C_{i,i+1}$ to be $|++\rangle\langle++|$ for $i>2$, and for remaining two-term-per-side inequality, we can derive the following:
$\text{CNOT}_{1,2}(B_{1,2}\otimes I_{3}+I_{1}\otimes B_{2,3})\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $\text{CNOT}_{1,2}B_{1,2}\text{CNOT}_{1,2}^{{\dagger}}\otimes I_{3}+\text{CNOT}_{1,2}I_{1}\otimes B_{2,3}\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $\text{CNOT}_{1,2}|0+\rangle\langle 0+|\text{CNOT}_{1,2}^{{\dagger}}\otimes I_{3}+\text{CNOT}_{1,2}I_{1}\otimes|++\rangle\langle++|\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+\text{CNOT}_{1,2}(|0\rangle\langle 0|+|1\rangle\langle 1|)\otimes|++\rangle\langle++|\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+|0\rangle\langle 0|\otimes|++\rangle\langle++|+\text{CNOT}_{1,2}|1\rangle\langle 1|\otimes|++\rangle\langle++|\text{CNOT}_{1,2}^{{\dagger}}\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+|0\rangle\langle 0|\otimes|++\rangle\langle++|+\text{CNOT}_{1,2}|1\rangle\langle 1|\otimes|+\rangle\langle+|\text{CNOT}_{1,2}^{{\dagger}}\otimes|+\rangle\langle+|\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
$\Longleftrightarrow$ $|0+\rangle\langle 0+|\otimes I_{3}+I_{1}\otimes|++\rangle\langle++|\leq C_{1,2}\otimes I_{3}+I_{1}\otimes C_{2,3}$
where, in the third-to-last and last steps, we use the following facts:
[…]
Therefore, we find that the post-state $\mathscr{C}$ predicate is
(18) $(C_{1,2},C_{2,3},C_{3,4}\cdots,C_{n-1,n})=(|0+\rangle\langle 0+|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|).$
(coinciding with the pre-state predicate $\mathscr{B}=(B_{1,2},B_{2,3},\dots,B_{n-1,n})$), and the post-state predicate is
$\mathcal{R}=(R_{1,2},R_{2,3},\cdots,R_{n-1,n})=(|00\rangle\langle 00|+|11\rangle\langle 11|,|00\rangle\langle 00|+|10\rangle\langle 10|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|).$
Equation˜18 illustrates an advantage of our choice of predicates. Because of properties such as those given in Equation˜17, Equation˜12 remains invariant under the application of CNOT gates. This invariance allows us to derive the strongest postcondition, while preserving the local structure of the matrix representation of predicates. As a result, we were able to make choices that made the inequalities that we worked with tight (or saturated, i.e., satified as equalities), making it easier to determine the postcondition. The right-hand side of Equation˜18 continues to serve as the predicate of local observables, as reasoning continues about the remaining CNOT gates.
After applying the CNOT gate on $q_{1}$ and $q_{r}$, the postcondition can be chosen as $(\mathscr{D}\mid\mathcal{S})$, where
$\mathscr{D}$ $=(|0+\rangle\langle 0+|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|)$
$\mathcal{S}$ $=(|00\rangle\langle 00|+|11\rangle\langle 11|,\cdots,|00\rangle\langle 00|+|11\rangle\langle 11|,|00\rangle\langle 00|+|10\rangle\langle 10|,|00\rangle\langle 00|,\cdots,|00\rangle\langle 00|).$
At the right end of the circuit, after the application of $U_{1}\otimes U_{2}\otimes\cdots\otimes U_{n}$—where each $U_{i}$ is a single-qubit unitary—the locality structure of the predicates remains unchanged. Therefore, we can choose the postcondition to be $(\mathscr{F}\mid\mathcal{T})$, where
$\mathscr{F}$ $=(\beta_{1}\otimes\beta_{2},\cdots,\beta_{n-1}\otimes\beta_{n})$
$\mathcal{T}$ $=(\psi_{1}\otimes\psi_{2}+\phi_{1}\otimes\phi_{2},\cdots,\psi_{n-1}\otimes\psi_{n}+\phi_{n-1}\otimes\phi_{n})$
with $\beta_{i}=|\beta_{i}\rangle\langle\beta_{i}|,\psi_{i}=|\psi_{i}\rangle\langle\psi_{i}|,\ \ \phi_{i}=|\phi_{i}\rangle\langle\phi_{i}|$ and
[…]
Let the output state be $\rho=|\Psi\rangle\langle\Psi|$. By ˜3.8, our proof of $\mathscr{F}$ implies that
(19) $\frac{n-1}{4}=\sum_{i=1}^{n-1}{\mathrm{Tr}}(|00\rangle\langle 00||++\rangle\langle++|)\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\beta_{i}\otimes\beta_{i+1})].$
STEP 2. This step mirrors Step 1, starting from the precondition $((|--\rangle\langle--|,|--\rangle\langle--|,\dots,|--\rangle\langle--|)\mid\mathcal{P})$. The postcondition of the full circuit can then be computed as
[…]
These conditions imply that
(20) $\frac{n-1}{4}\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\delta_{i}\otimes\delta_{i+1})].$
STEP 3 (QAI influence on QHL). According to Theorem 2.2, $\mathcal{Q}$ implies that the output state of the GHZ circuit is of form
[…]
for complex numbers $|a|^{2}+|b|^{2}=1$.
According to $\langle\psi_{i}|\phi_{i}\rangle=0$, we have $\rho_{i,i+1}=|a|^{2}\psi_{i}\otimes\psi_{i+1}+|b|^{2}\phi_{i}\otimes\phi_{i+1}.$ Then
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\beta_{1}\otimes\beta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall i>1$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\beta_{1}\otimes\beta_{2})]=0,\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall i>1$
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\delta_{1}\otimes\delta_{2})]=0,\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall i>1$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\delta_{1}\otimes\delta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall i>1.$
Equations˜19 and 20 imply
$\frac{n-1}{4}\leq|a|^{2}/2+\sum_{i=2}^{n-1}\frac{|a|^{2}+|b|^{2}}{4}=|a|^{2}/2+\frac{n-2}{4}\Longrightarrow\frac{1}{2}\leq|a|^{2}$
$\frac{n-1}{4}\leq|b|^{2}/2+\sum_{i=2}^{n-1}\frac{|a|^{2}+|b|^{2}}{4}=|b|^{2}/2++\frac{n-2}{4}\Longrightarrow\frac{1}{2}\leq|b|^{2}$
Together with $|a|^{2}+|b|^{2}=1$, we have $|a|^{2}=|b|^{2}=\frac{1}{2}$. In other words, there exists $\theta$ such that
[…]
Equation˜21 represents a closed-form expression for the circuit’s output with a single unknown real parameter. Such quantitative reasoning cannot be achieved with qualitative methods like QAI alone.
Moreover, the reasoning process scales with the number of qubits in the circuit: during the forward-reasoning process described above, for the reasoning steps carried out for each gate, the total size of the matrices that represent local observables and local projections in the pre- and post-conditions is always linear in the number of qubits. There are $O(n)$ gates; hence, the total amount of space needed to write down the SAQR-QC proof is $O(n^{2})$.
5. Quantum Phase Estimation
We apply our proof system to analyze Quantum Phase Estimation (QPE), a foundational algorithm central to Shor’s algorithm, whose intricate entanglement and nuanced success bounds have historically resisted scalable formal verification. Our approach unfolds in two stages: first, we establish the correctness of the Quantum Fourier Transform (QFT) via Quantum Abstract Interpretation (QAI); second, we leverage this foundation to analyze the full QPE circuit within our logic framework, deriving a lower bound on its success probability. To our knowledge, this constitutes the first scalable formal verification of QPE and QFT, offering a compositional, tractable methodology for reasoning about both correctness and probabilistic performance in key quantum algorithms.
5.1. Quantum Fourier transform
The Quantum Fourier Transform (QFT)—the quantum analogue of the discrete Fourier transform (see Figure˜4)—forms the computational core of landmark algorithms such as Shor’s factoring and quantum phase estimation. We introduce a lossless local reasoning framework for the QFT grounded in Quantum Abstract Interpretation (QAI). Our method derives an abstract output state whose concretization exactly coincides with the true QFT output for all computational-basis inputs. To the best of our knowledge, this constitutes the first verification framework that provides proofs polynomial in the size of the program while achieving such exact semantic correspondence for the QFT—and, by extension, for quantum phase estimation.
Figure 4. Quantum Fourier Transform: swap gates that reverse the qubit order at the circuit’s end are omitted.
The circuit uses Hadamard and phase gates $R_{m}$; for $m>2$, $R_{m}$ lies outside the Clifford group. See Subsection 2.4 for definitions of $H$ and $R_{m}$.
[…]
$\mathcal{P}=(P_{1},P_{2},\cdots,P_{n})=(|j_{1}\rangle\langle j_{1}|,|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|).$
In Appendix E, we prove the following using QAI
(23) $\vDash^{QAI}\{\mathcal{P}\}QFT\{(\psi_{j_{n}},\psi_{j_{n-1}j_{n}},\cdots,\psi_{j_{2}\cdots j_{n}},\psi_{j_{1}\cdots j_{n}})\}$
The postcondition derived from QAI is an abstract state represented as a tuple of density matrices corresponding to pure quantum states. By applying the concretization function from ˜2.2, we infer that the concrete state lies in the subspace
$\psi_{j_{n}}\otimes\psi_{j_{n-1}j_{n}}\otimes\cdots\otimes\psi_{j_{2}\cdots j_{n}}\otimes\psi_{j_{1}\cdots j_{n}}.$
Because the space so defined is a 1-dimensional subspace, it follows that the density matrix of the output state must exactly equal the pure-state projection onto this vector—that is, $\psi_{j_{n}}\otimes\psi_{j_{n-1}j_{n}}\otimes\cdots\otimes\psi_{j_{2}\cdots j_{n}}\otimes\psi_{j_{1}\cdots j_{n}}$.
During the forward-reasoning process described above, for the reasoning steps carried out for each gate, the total size of the matrices that represent local observables and local projections in the pre- and post-conditions is always linear in the number of qubits. There are $O(n^{2})$ gates; hence, the total amount of space needed to write down the proof is $O(n^{3})$.
5.2. Quantum Phase Estimation
In this section, we present a quantitative local analysis of the Quantum Phase Estimation (QPE) algorithm by combining both backward and forward reasoning techniques. We begin by decomposing the QPE algorithm into three constituent circuits and apply SAQR-QC to reason about each component individually—see Figure˜5. For an unknown phase $\theta$ and a given constant $k$, SAQR-QC can formally establish that—with probability at least $\frac{4}{\pi^{2}}$—the QPE algorithm produces an output bitstring whose last $k$ bits constitute the optimal $k$-bit approximation to the least significant $k$ bits of any $n$-bit binary representation/approximation of $\theta$.
[…]
We denote by $\mathbf{C_{1}}$ the segment of the program that precedes the application of the inverse Quantum Fourier Transform, $\mathcal{QFT}^{{\dagger}}$. The $\mathcal{QFT}^{{\dagger}}$ operation can be decomposed into two parts: the initial sequence of swap gates that reverses the order of the qubits, and the subsequent controlled rotation gates that implement the core of the inverse Fourier transform. We use $\mathbf{C_{2}}$ to refer specifically to the subcircuit following the swap gates within $\mathcal{QFT}^{{\dagger}}$.
Our analysis focuses on the least significant $k$ qubits, and proceeds in three steps. We will rely on the notation introduced in Equation˜30, and apply the QAI-based reasoning framework introduced in §˜3.6 to verify the behavior of $\mathbf{C_{1}}$.
Step 1: Reasoning about $\mathbf{C_{2}}$. Let $U$ be the circuit corresponding to $\mathbf{C_{2}}$. For $\mathbf{C_{2}}$, we choose the domain of the predicate as the first $k$ qubits of the first $n$ qubits, together with the last $m$ qubits—the qubits that $U$ acts upon. In Appendix F.1, we use backward reasoning to show
(24) $\{\mathscr{A}|\mathcal{I}\}\mathbf{C_{2}}\{(|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)|\mathcal{I}\}.$
[…]
Remark: Matrix multiplication will be efficient for constant dimension. Here, the dimension of matrices is $2^{k}$, which is a constant independent of $n$.
Step 2: Reasoning about SWAP gates. Now the precondition of Equation˜24, $\{\mathscr{A}|\mathcal{I}\}$, becomes the postconditon of the SWAP gates. Since the action of the SWAP gates only changes the last $k$ qubits into the first $k$ qubits, we may just write the precondition of the SWAP gates as $\{\mathscr{A}^{\prime}|\mathcal{I}\}$ with $\mathscr{A}^{\prime}=(|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|)$ on the first $k$ qubits and the last $m$ qubits. These arguments prove
(25) $\{\mathscr{A}^{\prime}|\mathcal{I}\}\mathbf{S}\{\mathscr{A}|\mathcal{I}\}$
where we use $\mathbf{S}$ to denote the SWAP gates. [^3] We change the abstract domain for simplicity of presentation. This change will not affect our statement’s correctness because the SWAP gates’ action is clear. If we want to fix the abstract domain, we can consider $(s_{1},\cdots,s_{m})$ with $s_{1}$ being the last $k$ qubits together with the last $m$ qubits; $s_{2}$ being the result of applying the first SWAP gate on $s_{1}$; $\cdots$; $s_{m}$ being the result of applying the last SWAP gate on $s_{m-1}$.
Step 3: Reasoning about $\mathbf{C_{1}}$. For $\mathbf{C_{1}}$, we select the predicate domain to include the first $k$ qubits of the initial $n$ and the last $m$ qubits where $U$ acts. In Appendix F.2, we apply our proof rules to demonstrate
(26) $\{r|0\cdots 0\rangle\langle 0\cdots 0|\otimes|\psi\rangle\langle\psi|)|(|0\cdots 0\rangle\langle 0\cdots 0|\otimes|\psi\rangle\langle\psi|\}\mathbf{C_{1}}\{\mathscr{A}^{\prime}|\mathcal{P}\}$
where
$r$ $=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]=\frac{\sin^{2}(2^{n}\theta\pi)}{4^{k}\sin^{2}[(2^{n-k}\theta-0.j_{n-k+1}\cdots j_{n})\pi]}$
$\mathcal{P}$ $=(|\omega\rangle\langle\omega|\otimes|\psi\rangle\langle\psi|)$
[…]
$r$ $=\frac{\sin^{2}(2^{n}\theta\pi)}{4^{k}\sin^{2}[(2^{n-k}\theta-0.j_{n-k+1}\cdots j_{n})\pi]}=\frac{\sin^{2}(2^{n}\epsilon\pi)}{4^{k}\sin^{2}(2^{n-k}\epsilon\pi)}\geq\frac{|2\cdot 2^{n}\epsilon|^{2}}{4^{k}\cdot|2^{n-k}\epsilon\cdot\pi|^{2}}\geq\frac{4}{\pi^{2}}.$
[…]
6. Related Work
Combining Logic and Abstract Interpretation. Combining Hoare logic with abstract interpretation is a well-established strategy in classical program analysis. Many software model-checking tools integrate logic-driven state-space exploration with dataflow information obtained via abstract interpretation, often computed as a preliminary pass. A prominent example is the Seahorn solver for Constrained Horn Clauses, whose Crab component provides abstract-interpretation capabilities (Crab, ; GN21, ). This integration illustrates how abstract interpretation can complement logical reasoning, resulting in verification that is both scalable and precise. Our work extends this principle to the quantum domain, where the combination supports reasoning about program semantics that would be difficult to achieve with either technique in isolation.
Quantum Hoare Logic. Researchers have developed a variety of quantum Hoare logics. A comparative study in (RA17, ) examined three representative logics (CMS, ; 10.1007/978-3-642-10622-4_7, ; Ying11, ). Quantum Hoare logics can be broadly categorized into expectation-based and satisfaction-based approaches. Following the seminal work of D’Hondt and Panangaden (DP06, ), expectation-based approaches (Ying11, ; BHY19, ; li2019quantum, ; feng2021quantum, ) use positive operators as assertions for quantum states, defining the expectation that a state $\rho$ satisfies an assertion $M$ as ${\mathrm{Tr}}(M\rho)$. In contrast, satisfaction-based logics—exemplified by Zhou et al. (ZYY19, ) and indepedently by Unruh (Unruh, ; Unruh2, )—treat subspaces of the Hilbert space as assertions, providing a direct semantic correspondence between program correctness and the geometry of quantum states. Here, a state $\rho$ satisfies an assertion $P$ if the support of $\rho$ is included in $P$. In all of these works, predicates $M$ or $P$ are represented as $2^{n}\times 2^{n}$ matrices, which precludes efficient computation and limits scalability for larger quantum systems.
Quantum Separation logic. Quantum reasoning frameworks have made significant progress along multiple fronts. Quantum separation logic, introduced in (zhou2021quantum, ), extends the principles of Bunched Implications (o'hearn_pym_1999, ) to the quantum setting, enabling local reasoning about quantum programs. The framework in (10.1145/3498697, ) further incorporates hybrid state spaces by supporting both classical variables and the dynamic allocation and deallocation of quantum qubits. However, in both approaches, the separating conjunction is interpreted strictly as a tensor product—corresponding to quantum independence—which substantially limits the expressiveness and applicability of the logic. In addition, because predicates are restricted to projection operators, these systems support only qualitative rather than quantitative reasoning about quantum states and their interactions.
Quantum abstract interpretation. (YP21, ) presented an approach to quantum abstract interpretation for reasoning about quantum circuits, using the satisfaction-based approach. Other works, such as (Bichsel_2023, ), investigate the abstract interpretation of quantum programming using variants of the Gottesman-Knill theorem.
Symbolic abstraction, strongest consequence, and weakest sufficient condition. The inexpressibility issues discussed in §˜3.2 are a manifestation of the constraints that one faces when working with an “impoverished” logic (or logic fragment). These issues have been studied in the context of abstract interpretation as what is (now) called the symbolic-abstraction problem (DBLP:conf/vmcai/RepsSY04, ; DBLP:conf/cav/ThakurR12, ), and phrased in purely logical terms as the strongest-consequence problem (DBLP:conf/vmcai/RepsT16, , §5), as follows: Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the strongest formula $\psi\in\mathcal{L}^{\prime}$ such that $\varphi\vDash\psi$. The strongest-consequence problem naturally arises in approximate forwards reasoning, to over-approximate a postcondition. The discussion in §˜3.2 concerned backwards reasoning for which one faces the dual problem, the weakest sufficient-condition problem: Given formula $\varphi\in\mathcal{L}$, and a different logic $\mathcal{L}^{\prime}$, find the weakest formula $\chi\in\mathcal{L}^{\prime}$ such that $\chi\vDash\varphi$.
As we saw in §˜3.2, the strongest consequence or weakest sufficient condition may not always be expressible in $\mathcal{L}^{\prime}$, in which case one has to fall back on finding what one hopes is a suitably strong consequence or a suitably weak sufficient condition, respectively. Scherpelz et al. (DBLP:conf/pldi/ScherpelzLC07, ) presented a best-effort method for computing sufficient conditions as part of an algorithm for creating abstract transformers for use with parameterized predicate abstraction (DBLP:conf/birthday/Cousot03, ). Their method performs weakest liberal precondtion (WLP) of a post-state predicate with respect to a concrete transformer $\tau$, and then uses heuristics to identify combinations of pre-state predicates that entail the WLP value.
Reasoning about Shor’s factoring. The QFT circuit contains many non-Clifford gates, so variants of the Gottesman-Knill theorem do not directly apply. Shor’s algorithm has been formally verified in (feng2021quantum, ; https://doi.org/10.48550/arxiv.2204.07112, ). The former uses an expectation-based approach with classical variables, while the latter employs the Coq proof assistant and the Small Quantum Intermediate Representation (sQIR) (Hicks21, ). Since Coq operates symbolically, the reasoning about quantum phase estimation in (https://doi.org/10.48550/arxiv.2204.07112, ) is symbolic.
Symbolic verification. There has been work to extend symbolic-verification techniques to the quantum domain (10.1145/3725725, ; CACM:CCLLTY25, ), using logic- and automata-based techniques developed for symbolic verification of classical programs to analyze the correctness of quantum programs. In contrast, SAQR-QC is not symbolic in nature. Reasoning steps in SAQR-QC can involve matrix multiplications and other mathematical operations on specific values of specific sizes.
SAQR-QC also does not support parameterized reasoning about families of circuits parameterized on the number of qubits. On the contrary, SAQR-QC can be used to reason about a specific circuit with a specific number of qubits.
Packing of variables in abstract domains. The motivation for using local observables and local projections defined with respect to a tuple $(s_{1},\cdots,s_{m})$ of sets of qubit indexes is to make reasoning scalable: each reasoning step involves only a small number of qubits. This idea is similar to the idea of “packing” variables in numeric domains, as used in Astrée (DBLP:conf/pldi/BlanchetCCFMMMR03, ): a program’s numeric variables are placed in sets (“packs”) so that each abstract transformer of a numeric abstract domain can operate on a single pack at a time. As with our qubit sets, each variable can be in multiple packs.
7. Conclusion and Future work
This paper introduces a framework for local reasoning about quantum circuits that is both scalable and addresses quantitative properties. The framework combines Quantum Hoare Logic, which tracks quantitative properties of quantum states, with quantum abstract interpretation, which provides scalable reasoning about qualitative properties. Unlike a prior scalable framework that is restricted to Clifford circuits (GottesmanKnill, ), our approach supports non-Clifford gates, greatly extending its applicability. By tracking tuples of reduced density matrices via tuples of local observables, it fills a long-standing gap in methods for reasoning about quantum programs, enabling scalable and quantitative verification of quantum circuits on classical computers.
We apply our logic to GHZ circuits with non-Clifford unitaries, successfully characterizing the output state up to a relative phase. We further demonstrate its use on the Quantum Fourier Transform, and derive a lower bound on the success probability of the Quantum Phase Estimation algorithm.
Extending our methods to general quantum programs with classical control (Ying11, ; zhou2021quantum, ) and designing a logic for noisy programs (10.1145/3453483.3454029, ; Wu19, ; ZYY19, ) are promising directions for future work. Another direction, automating our reasoning framework, will require solving two fundamental challenges: (i) how to select qubit domains, and (ii) how to synthesize predicates as positive semidefinite observables under semidefinite constraints.
Appendix A Appendix A: Lemmas and Proofs of Lemmas
Lemma A.1.
$P\subseteq Q\Longrightarrow PQ=QP=P.$
Lemma A.2.
$A\leq B\Longrightarrow PAP\leq PBP.$
˜2.1. Let $\rho$ be the density matrix of an $n$-qubit system, and let $s\subseteq[n]$. Then for any observable $A_{s}$ acting on subsystem $s$,
${\mathrm{Tr}}\left((A_{s}\otimes I_{[n]\setminus s})\rho\right)={\mathrm{Tr}}\left(A_{s}\rho_{s}\right).$
Proof.
We express $\rho$ as
$|,\rho=\sum_{i,j}\rho_{i,j}\otimes\ket{i}\bra{j}$,
[…]
$=\sum_{i}{\mathrm{Tr}}(A_{s}\rho_{i,i})$
$={\mathrm{Tr}}\left(A_{s}\sum_{i}\rho_{i,i}\right)$
$={\mathrm{Tr}}\left(A_{s}\rho_{s}\right),$
where $\rho_{s}={\mathrm{Tr}}_{[n]\setminus s}(\rho)=\sum_{i}\rho_{i,i}$.
Appendix B Appendix C: Proof of Theorem 3.1
Proof.
Assume that $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ as defined in ˜3.8. Then for any $\rho$, we have
${\mathrm{Tr}}(M_{\mathscr{P}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$
Let us choose $\rho\models\mathcal{P}$. According to Lemma 2.5, we have
${\mathrm{Tr}}(M_{\mathscr{P}}\rho)={\mathrm{Tr}}(\sum_{i}P_{s_{i}}\rho_{s_{i}})=\sum_{i}{\mathrm{Tr}}(P_{s_{i}}\rho_{s_{i}})=\sum_{i}1=m.$
On the other hand,
${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$
$=$ ${\mathrm{Tr}}\left(\left(\sum_{i}Q_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\llbracket\mathbf{C}\rrbracket(\rho)\right)$
$=$ $\sum_{i}{\mathrm{Tr}}(Q_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))$
$\leq$ $\sum_{i}{\mathrm{Tr}}(I_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))$
$=$ $\sum_{i}1=m.$
where the inequality follows from ˜2.4 and $Q_{s_{i}}\leq I_{s_{i}}$.
Therefore, ${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$ and ${\mathrm{Tr}}(Q_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))=1$ for all ii. According to Lemma 2.5, we know ${\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho)\vDash Q_{s_{i}}$, which implies
$\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\},$
where LABEL:qai $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$ iff for all $\rho$, $\rho\vDash\mathcal{P}$ implies $\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q}$. ∎
Appendix C Appendix C: Proof of Theorem 3.2
Proof.
We prove the soundness of each rule.
Rule Skip
$\textsc{Skip}\quad\ \frac{}{\{\mathscr{A}|\mathcal{P}\}\mathbf{Skip}\{\mathscr{A}|\mathcal{P}\}}$
For any $\rho\vDash^{QAI}\mathcal{P}$, we have $\llbracket\mathbf{Skip}\rrbracket(\rho)=\rho\vDash^{QAI}\mathcal{P}$. Moreover,
${\mathrm{Tr}}[M_{\mathscr{A}}(\rho)]\leq{\mathrm{Tr}}[M_{\mathscr{A}}(\llbracket\mathbf{Skip}\rrbracket(\rho))].$
Rule Unit
$\textsc{Unit}\quad\ \frac{\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})}{\{\mathscr{A}|\mathcal{P}\}{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\{\mathscr{B}|U_{F}^{\sharp}(\mathcal{P})\}}$
For $\rho\vDash^{QAI}\mathcal{P}$, using Theorem 2.1, we have
$\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\rho)=U_{F}\rho U_{F}^{{\dagger}}\vDash^{QAI}U_{F}^{\sharp}(\mathcal{P}).$
Moreover, we have $\rho\vDash\gamma(\mathcal{P})$; that is, $\rho=\gamma(\mathcal{P})\rho\gamma(\mathcal{P})$. According to Lemma 2.2, we know
${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={{\mathrm{Tr}}}(M_{\mathscr{A}}\gamma(\mathcal{P})\rho\gamma(\mathcal{P}))={\mathrm{Tr}}(\rho\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P}))$
On the other hand, Lemma 2.2 also implies
${{\mathrm{Tr}}}(M_{\mathscr{B}}\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\rho))$
$=$ ${{\mathrm{Tr}}}[M_{\mathscr{B}}\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\gamma(\mathcal{P})\rho\gamma(\mathcal{P})]$
$=$ ${\mathrm{Tr}}[M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})\rho\gamma(\mathcal{P})U_{F}^{{\dagger}}]$
$=$ ${\mathrm{Tr}}[\rho\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})].$
Therefore, the condition
$\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P})\leq\gamma(\mathcal{P})\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\gamma(\mathcal{P})$
implies that, for $\rho\vDash^{QAI}\mathcal{P}$, we have
${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}(\rho\gamma(\mathcal{P})M_{\mathscr{A}}\gamma(\mathcal{P}))\leq{\mathrm{Tr}}[\rho\gamma(\mathcal{P})U_{F}^{{\dagger}}M_{\mathscr{B}}U_{F}\gamma(\mathcal{P})]\leq{{\mathrm{Tr}}}(M_{\mathscr{B}}\llbracket{\overline{q}}:=U_{F}\left[{\overline{q}}\right]\rrbracket(\rho))$
by invoking Lemma 2.2.
This argument proves Rule Unit.
Rule Seq
$\textsc{Seq}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1}\{\mathscr{D}|\mathcal{R}\}\ \ \ \ \ \ \{\mathscr{D}|\mathcal{R}\}\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1};\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}}$
For $\rho\vDash^{QAI}\mathcal{P}$, we have
$\{\mathscr{A}|\mathcal{P}\}\mathbf{C}_{1}\{\mathscr{D}|\mathcal{R}\}\Longrightarrow\llbracket\mathbf{C}_{1}\rrbracket(\rho)\vDash^{QAI}\mathcal{R},\ \ \ {\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{D}}(\llbracket\mathbf{C}_{1}\rrbracket(\rho))].$
According to $\{\mathscr{D}|\mathcal{R}\}\mathbf{C}_{2}\{\mathscr{B}|\mathcal{Q}\}$ and $\llbracket\mathbf{C}_{1}\rrbracket(\rho)\vDash^{QAI}\mathcal{R}$, we obtain that
$\llbracket\mathbf{C}_{2}\rrbracket(\llbracket\mathbf{C}_{1}\rrbracket(\rho))\vDash^{QAI}\mathcal{Q}$
${\mathrm{Tr}}[M_{\mathscr{D}}(\llbracket\mathbf{C}_{1}\rrbracket(\rho))]\leq{\mathrm{Tr}}[\llbracket M_{\mathscr{B}}(\mathbf{C}_{2}\rrbracket(\llbracket\mathbf{C}_{1}\rrbracket(\rho)))]$
According to the fact that $\llbracket\mathbf{C}_{1};\mathbf{C}_{2}\rrbracket(\rho))=\llbracket\mathbf{C}_{2}\rrbracket(\llbracket\mathbf{C}_{1}\rrbracket(\rho))$, we know that if $\rho\vDash^{QAI}\mathcal{P}$,
$\llbracket\mathbf{C}_{1};\mathbf{C}_{2}\rrbracket(\rho))\vDash^{QAI}\mathcal{Q}$
${\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{B}}(\llbracket\mathbf{C}_{1};\mathbf{C}_{2}\rrbracket(\rho))]$
This argument proves Rule Seq.
Rule Con
$\textsc{Con}\quad\frac{\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\},\ \ \mathscr{D}\sqsubseteq\mathscr{A},\ \mathscr{B}\sqsubseteq\mathscr{E},\ \mathcal{R}\sqsubseteq\mathcal{P},\ \mathcal{Q}\sqsubseteq\mathcal{T}}{\{\mathscr{D}|\mathcal{R}\}\mathbf{C}\{\mathscr{E}|\mathcal{T}\}}$
For any $\rho\vDash^{QAI}\mathcal{R}$, we have
$\rho\vDash^{QAI}\mathcal{R}\sqsubseteq\mathcal{P}.$
Moreover,
$\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}\Longrightarrow\llbracket\mathbf{C}\rrbracket(\rho)\vDash^{QAI}\mathcal{Q}\sqsubseteq\mathcal{T},\ \ {\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{B}}(\llbracket\mathbf{C}\rrbracket(\rho))].$
According to Lemma 3.1, we have
$M_{\mathscr{D}}\leq M_{\mathscr{A}},\ \ M_{\mathscr{B}}\leq M_{\mathscr{E}}$
$\Longrightarrow$ ${\mathrm{Tr}}[M_{\mathscr{D}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{A}}\rho]\leq{\mathrm{Tr}}[M_{\mathscr{B}}(\llbracket\mathbf{C}\rrbracket(\rho))]\leq{\mathrm{Tr}}[M_{\mathscr{E}}(\llbracket\mathbf{C}\rrbracket(\rho))].$
This argument proves the correctness of Rule Con. ∎
Appendix D Appendix D: Equation (16) implies Equation (11)
(27) for all $j$ $\gamma(\mathcal{P})P_{j}\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)P_{j}\gamma(\mathcal{P})$ $\leq\gamma(\mathcal{P})P_{j}U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)UP_{j}\gamma(\mathcal{P})$
(28) $\Longrightarrow$ for all $j$ $\gamma(\mathcal{P})\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\gamma(\mathcal{P})$ $\leq\gamma(\mathcal{P})U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\gamma(\mathcal{P})$
$\Longrightarrow$ $\sum_{j}\gamma(\mathcal{P})\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\gamma(\mathcal{P})$ $\leq\sum_{j}\gamma(\mathcal{P})U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\gamma(\mathcal{P})$
(29) $\Longrightarrow$ $\gamma(\mathcal{P})\left(\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)\gamma(\mathcal{P})$ $\leq\gamma(\mathcal{P})U^{{\dagger}}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\gamma(\mathcal{P})$
where in going from Equation˜16 to Equation˜27, we used ˜A.2; and in going from Equation˜27 to Equation˜28, we used Equation˜15 and ˜A.1.
Appendix E Appendix E: Reasoning about QFT using QAI
The quantum gates used in the circuit are the Hadamard gate and the phase gate $R_{m}$, $R_{m}$ does not belong to the Clifford group for $m>2$.
${H={\frac{1}{\sqrt{2}}}{\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}}\qquad{\text{and}}\qquad R_{m}={\begin{pmatrix}1&0\\ 0&e^{2\pi i/2^{m}}\end{pmatrix}}}$
[…]
We choose the domain $(\{1\},\{2\},\cdots,\{n\})$, and the precondition to be
$\mathcal{P}=(P_{1},P_{2},\cdots,P_{n})=(|j_{1}\rangle\langle j_{1}|,|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|).$
At the beginning of the program, according to $\beta_{k}=|j_{k}\rangle\langle j_{k}|\models_{p}P_{k}$, we have the input
[…]
After applying the first $H$ gate on $q_{1}$, we compute the postcondition, which becomes
$(H|j_{1}\rangle\langle j_{1}|H,|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)=(\psi_{j_{1}},|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
[…]
Applying the controlled-$R_{2}$ gate, we compute
$P_{1,2}:=\psi_{j_{1}}\otimes I_{2}\cap I_{1}\otimes|j_{2}\rangle\langle j_{2}|=\psi_{j_{1}}\otimes|j_{2}\rangle\langle j_{2}|.$
The postcondition becomes
$(\operatorname{supp}({\mathrm{Tr}}_{2}CR_{2}A_{1,2}CR_{2}^{{\dagger}}),\operatorname{supp}({\mathrm{Tr}}_{1}CR_{2}A_{1,2}CR_{2}^{{\dagger}}),|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
$=$ $(\psi_{j_{1}j_{2}},|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
[…]
$(\psi_{j_{1}j_{2}\cdots j_{n}},|j_{2}\rangle\langle j_{2}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
Next, we perform a similar procedure on the second qubit. The Hadamard gate puts us in the postcondition
$(\psi_{j_{1}j_{2}\cdots j_{n}},H|j_{2}\rangle\langle j_{2}|H,|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)=(\psi_{j_{1}j_{2}\cdots j_{n}},\psi_{j_{2}},|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
The controlled-$R_{2}$ through $R_{n-1}$ gates yield the predicate
$(\psi_{j_{1}j_{2}\cdots j_{n}},\psi_{j_{2}\cdots j_{n}},|j_{3}\rangle\langle j_{3}|,\cdots,|j_{n}\rangle\langle j_{n}|)$
We continue in this fashion for each qubit, giving a final predicate
$(\psi_{j_{1}j_{2}\cdots j_{n}},\psi_{j_{2}\cdots j_{n}},\psi_{j_{3}\cdots j_{n}},\cdots,\psi_{j_{n}}).$
It follows the Swap operation between the qubit $i$ and the qubit n+1in+1-i for $1\leq i\leq n$. After applying $SWAP(1,n)$, we obtain
$(\psi_{j_{n}},\psi_{j_{2}\cdots j_{n}},\cdots,\psi_{j_{n-1}j_{n}},\psi_{j_{1}\cdots j_{n}})$
by observing
$\operatorname{supp}({\mathrm{Tr}}_{n}SWAP(1,n)[\psi_{j_{1}\cdots j_{n}}\otimes I_{n}]\cap[I_{1}\otimes\psi_{j_{n}}]SWAP(1,n))$ $=\psi_{j_{n}}$
$\operatorname{supp}({\mathrm{Tr}}_{1}SWAP(1,n)[\psi_{j_{1}\cdots j_{n}}\otimes I_{n}]\cap[I_{1}\otimes\psi_{j_{n}}]SWAP(1,n))$ $=\psi_{j_{1}\cdots j_{n}}.$
After all the swap operations, the postcondition is
$(\psi_{j_{n}},\psi_{j_{n-1}j_{n}},\cdots,\psi_{j_{2}\cdots j_{n}},\psi_{j_{1}\cdots j_{n}})$
The postcondition derived from QAI is an abstract state represented as a tuple of density matrices corresponding to pure quantum states.
Appendix F Appendix F: Proofs of Eq. (24) and Eq. (26) in Section 5
F.1. Proof of Equation˜24
Equation˜24 is equivalent to
$M_{\mathscr{A}}\leq$ $U^{{\dagger}}[I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|\otimes|\psi\rangle\langle\psi|]U$
$=$ $U^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)U\otimes|\psi\rangle\langle\psi|.$
To see this, we observe
$H(1)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)H(1)^{{\dagger}}$
$=$ $CR_{n}(n,1)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)CR_{2}(n,1)^{{\dagger}}$
$=$ $CR_{2}(2,1)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)CR_{2}(2,1)^{{\dagger}}$
$=$ $H(n-k+2)(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)H(n-k+2)^{{\dagger}}$
$=$ $I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|.$
Let $V=H(1)^{-1}CR_{2}(2,1)^{-1}CR_{3}(3,1)^{-1}\cdots H(n-k+2)^{-1}$ means
$V^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)V=I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|$
We let $U=VU_{k}$ with $U_{k}$ being the sub-circuit which only applied on the last $k$-qubits, i.e.,
$U_{k}=H(n-k+1)^{-1}\cdots H(n-1)^{-1}CR_{2}(n,n-1)^{-1}H(n)^{-1},$
Then
$U^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)U\otimes|\psi\rangle\langle\psi|$
$=$ $U_{k}^{{\dagger}}(I_{1,\cdots,n-k}\otimes|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|)U_{k}\otimes|\psi\rangle\langle\psi|$
$=$ $I_{1,\cdots,n-k}\otimes U_{k}^{{\dagger}}|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|U_{k}\otimes|\psi\rangle\langle\psi|$
$U_{k}^{{\dagger}}$ is the standard Quantum Fourier transform applied on input $|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|$. Performing direct matrix multiplication (or according to the last subsection of reasoning about the Quantum Fourier transform), we know that
$U_{k}^{{\dagger}}|j_{n-k}\cdots j_{n}\rangle\langle j_{n-k}\cdots j_{n}|U_{k}=|\tau\rangle\langle\tau|$
This argument proves Equation˜24.
We first use quantum abstract interpretation (YP21, ) to prove
(31) $\{\cdot|(|0\cdots 0\rangle\langle 0\cdots 0|\otimes|\psi\rangle\langle\psi|)\}\mathbf{C_{1}}\{\cdot|\mathcal{P}\}.$
After applying the first $H$ gates, the post-condition becomes
$\{\cdot|(|+\cdots+\rangle\langle+\cdots+|\otimes|\psi\rangle\langle\psi|\}.$
For $CU^{2^{n-1}}\cdots CU^{2^{n-k}}$, direct matrix computation leads to the post-condition $\{\cdot|\mathcal{P}\}$.
For $CU^{2^{n-k+1}}\cdots CU^{2^{0}}$, there will be no change to the observable on the first $k$ qubits. For example, for the first gate $CU^{2^{0}}$, which is $CU$, we have the post-condition $\{\cdot||+\cdots+\rangle\langle+\cdots+|\otimes|\psi\rangle\langle\psi|\}$ by observing
$\operatorname{supp}[{\mathrm{Tr}}_{n}CU(|\omega\rangle\langle\omega|\otimes I_{n}\otimes|\psi\rangle\langle\psi|)CU^{{\dagger}}]$
$=$ $\operatorname{supp}[{\mathrm{Tr}}_{n}CU(|\omega\rangle\langle\omega|\otimes(|0\rangle\langle 0|+|1\rangle\langle 1|)\otimes|\psi\rangle\langle\psi|)CU^{{\dagger}}]$
$=$ $\operatorname{supp}[2|\omega\rangle\langle\omega|\otimes|\psi\rangle\langle\psi|]$
$=$ $|\omega\rangle\langle\omega|\otimes|\psi\rangle\langle\psi|.$
Similarly for $CU^{2^{n-k+1}}\cdots CU^{2^{0}}$. This argument proves Equation˜31.
F.2. Proof of Equation˜26
To prove Equation˜26, we observe
$|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|(r|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|)|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|$
$=$ $r|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|;$
and
$|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|[U(C_{1})^{{\dagger}}(|\tau\rangle\langle\tau|\otimes I\otimes|\psi\rangle\langle\psi|)U(C_{1})]|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|$
$=$ $|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|[V(C_{1})^{{\dagger}}(|\tau\rangle\langle\tau|\otimes I\otimes|\psi\rangle\langle\psi|)V(C_{1})]|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|$
$=$ $x|0\cdots 0\rangle\langle 0\cdots 0|\otimes I\otimes|\psi\rangle\langle\psi|;$
where
$V(C_{1})$ $=CU^{2^{n-k}}\cdots CU^{2^{n-1}}$
[…]
$={\mathrm{Tr}}[(|\tau\psi\rangle\langle\tau\psi|)V(C_{1})|0\cdots 0\psi\rangle\langle 0\cdots 0\psi|V(C_{1})^{{\dagger}}]$
$={\mathrm{Tr}}[|\tau\psi\rangle\langle\tau\psi||\omega\psi\rangle\langle\omega\psi|]$
$=|\langle\tau|\omega\rangle|^{2}$
[…]
$=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]$
$=r.$